### v1

# Unbiased Regression-Adjusted Estimation of Average Treatment Effects in Randomized Controlled Trials

Abstract
This article introduces a leave-one-out regression adjustment estimator (LOORA) for estimating average treatment effects in randomized controlled trials. The method removes the finite-sample bias of conventional regression adjustment and provides exact variance expressions for LOORA versions of the Horvitz-Thompson and difference-in-means estimators under simple and complete random assignment. Ridge regularization limits the influence of high-leverage observations, improving stability and precision in small samples. In large samples, LOORA attains the asymptotic efficiency of regression-adjusted estimator as in [lin2013agnostic], while remaining exactly unbiased. To construct confidence intervals, we rely on asymptotic variance estimates that treat the estimator as a two-step procedure, accounting for both the regression adjustment and the random assignment stages. Two within-subject experimental applications that provide realistic joint distributions of potential outcomes as ground truth show that LOORA eliminates substantial biases and achieves close-to-nominal confidence interval coverage.
leave-one-out estimators,
keywords:
[…]
[Name Surname; will be inserted later]
1 Introduction
Establishing causal relations is a central goal in many sciences, as well as economics and the social sciences. With observational data, selection bias often obscures the true effect of interventions. Randomized controlled trials (RCTs) overcome this problem by assigning treatments randomly, eliminating systematic differences between treatment and control groups, and enabling unbiased comparisons of outcomes. RCTs now stand at the core of modern empirical research in causal inference.
As experimental practice has matured, attention has shifted toward improving efficiency and statistical power. Researchers now collect extensive pretreatment information to predict outcomes and use regression adjustment to explain part of the outcome variance. This approach reduces noise and increases the precision of treatment effect estimation without requiring complex assignment mechanisms.
Most theoretical guarantees for regression-adjusted estimators rely on asymptotic arguments. However, early-stage clinical trials, marketing geo-experiments, and studies of firm- or country-level policies frequently operate in small sample regimes, where asymptotic results do not apply. In these cases, standard regression-adjusted estimators become biased, with undesirable consequences for treatment effect estimation and policy design. Moreover, high-leverage observations (that is, observations whose covariates strongly influence the regression fit) further degrade the performance of regression adjustment in small samples. This problem is well documented in influential work by freedman2008regression and young2019channeling.
To address these challenges, in this paper we develop a regression adjustment framework that is unbiased in finite samples and robust to influential observations. A key component of our analysis is a leave-one-out regression adjustment (LOORA) procedure, which removes the bias of classical regression adjustment. Within this framework, we develop two estimators of the average treatment effect (ATE): LOORA-HT, a Horvitz–Thompson estimator for simple random assignment, and LOORA-DM, a difference-in-means estimator for complete random assignment. LOORA-HT and LOORA-DM are easy to implement and deliver a performance that previously required complex assignment mechanisms (see harshaw2022design).
LOORA-HT and LOORA-DM simultaneously address three core problems in regression-adjusted estimation of treatment effects. First, they answer the critique of freedman_2008_b by providing unbiased, randomization-based estimators. Second, they reach the asymptotic efficiency of lin2013agnostic. Third, by stabilizing estimation in the presence of influential observations, they resolve the issues identified by young2019channeling.
In two within-subject experiments that provide a realistic ground truth for the joint distribution of potential outcomes, LOORA-HT and LOORA-DM substantially reduce estimation bias and yield confidence intervals with close-to-nominal coverage.
1.1 Related Work
The practice of regression adjustment in randomized experiments dates at least to the use of analysis of covariance (ANCOVA) in classical experimental design, the effects of pre-treatment covariates are removed through linear adjustment in order to improve precision. [^1] See, for example, discussions in fisher1971design. Despite its pervasiveness, the formal large-sample properties of regression adjustment after randomization have been repeatedly questioned. In two influential critiques, freedman2008regression; freedman_2008_b argue that ordinary least squares (OLS) adjustment can introduce bias (in both point estimates and standard errors) when the linear working model is misspecified, and can perform worse than the unadjusted difference-in-means estimator in finite samples.
In response, lin2013agnostic shows that, under complete random assignment, an OLS regression of outcomes on treatment, covariates, and all treatment-by-covariate interactions delivers an estimator that is consistent for the average treatment effect and asymptotically is at least as efficient as the unadjusted difference-in-means estimator, even when the linear model is misspecified. These results clarify the large sample properties of regression adjustment after randomization. However, two gaps remain in the literature: (i) lin2013agnostic does not provide a design-based finite-sample unbiasedness guarantee for regression adjustment, and (ii) the interacted specification in lin2013agnostic can be unstable in moderate samples because it increases the number of regressors, adding with additional interaction terms and potentially inflating leverage scores and sensitivity to influential units.
Our paper contributes to this literature by constructing estimators—LOORA-HT under simple random assignment and LOORA-DM under complete random assignment—that (i) are exactly unbiased for the finite-population average treatment effect, (ii) admit closed-form variance expressions in finite samples, and (iii) are asymptotically efficient, attaining the variance bound of lin2013agnostic. A key technical component is leave-one-out regression adjustment, in which each unit’s outcome is adjusted using coefficients estimated from a regression that excludes that unit. This leave-one-out procedure removes the finite-sample bias identified by freedman2008regression; freedman_2008_b. In addition, LOORA-HT and LOORA-DM use of ridge regression to control the impact of high-leverage observations. For the difference-in-means setting, we show that LOORA-DM admits a representation that is equivalent to a leave-two-out construction in the sense of spiess2025optimal, thereby satisfying necessary conditions for unbiasedness under complete random assignment. In this way, our estimators provide constructive finite-sample analogues of the asymptotic guarantees in lin2013agnostic, while remaining valid without requiring parametric outcome models.
A related line of work examines precision gains in randomized studies through experimental design rather than regression adjustment. harshaw2022design propose a Gram–Schmidt walk design, which chooses treatment assignments in a way that approximately balances covariates and, in turn, controls the variance of the Horvitz–Thompson estimator through ridge-regularization. Their goal is to optimize the randomization design to deliver low-variance Horvitz–Thompson estimators in finite samples.
In contrast, we keep the assignment mechanism fixed (simple or complete random assignment) and instead adjust outcomes through leave-one-out regressions. This distinction matters in practice, because our methods can be applied to commonly used experimental designs without the need to change the assignment rule.
ghadiri2023finite introduce a cross-fitted regression-adjusted Horvitz–Thompson estimator under simple random assignment. Their estimator splits the sample into two folds, estimates regression coefficients on one fold, and applies those coefficients to adjust outcomes in the other fold, yielding an exactly unbiased estimator with an explicit nonasymptotic variance bound. However, the bound in ghadiri2023finite is looser than the lin2013agnostic asymptotic efficiency bound (by roughly a constant factor) and they do not develop confidence intervals.
Our leave-one-out construction both tightens the variance bound and delivers an estimator whose asymptotic variance matches lin2013agnostic. We also provide a simple two-step regression view that yields a Huber–White style variance estimator and confidence intervals that nearly match the nominal coverage values on two applications.
lei2021regression propose a bias-correction procedure for regression adjustment when the number of covariates grows with the sample size. Their method removes higher-order bias terms but does not yield exact unbiasedness in finite samples. chang2024exact develop an exact bias-correction approach that eliminates bias entirely, although they do not provide a finite-population variance bound for the resulting estimator. armstrong2021finite study the construction of optimal confidence intervals in finite samples under the assumption that regression errors are normally distributed with known variances. Finally, kline2020leave analyze leave-one-out methods for variance estimation in linear models, providing asymptotic results but not nonasymptotic guarantees.
A large concurrent literature studies regression adjustment and inference in more structured designs: stratified or blocked experiments, matched pairs, cluster randomization, or covariate-adaptive randomization. For example, bai2024primer survey modern design-based analysis of randomized experiments, emphasizing the role of stratification, regression adjustment and cluster randomized experiments. bai2025new; bai2025inference develop design-based variance estimators and inference procedures for finely stratified and matched-pair experiments, including settings with imperfect compliance. cytrynbaum2024covariate characterizes, for general stratified randomization schemes, the asymptotically optimal linear covariate adjustment and shows that the classical interacted regression of lin2013agnostic can be inefficient away from complete random assignment. We also note that stratification becomes impractical in the presence of high-dimensional covariates, whereas our regression-adjustment approach remains applicable in such settings.
1.2 Notation
Vectors and matrices.
We denote matrices and vectors by bold uppercase and lowercase letters, respectively. The $i^{\text{th}}$ entry of a vector $\bm{u}$} is denoted by $u_{i}$. The transposed $i^{\text{th}}$ row of a matrix $\bm{X}$} is denoted by ${\bm{x}}_{i}$, and its $(i,j)^{\text{th}}$ entry by $x_{ij}$. For a constant $c$ and a vector $\bm{u}$}, expressions such as $c+{\bm{u}}$ and ${\bm{u}}^{-1}$ are interpreted entrywise. When $\bm{u}$} and $\bm{v}$} are vectors of equal dimension, ${\bm{u}}{\bm{v}}$ and ${\bm{u}}/{\bm{v}}$ denote entrywise (Hadamard) product and division, respectively. A vector’s associated diagonal matrix appears in uppercase; for instance, $\boldsymbol{T}$ denotes the diagonal matrix $\bm{t}$} as its main diagonal. For any vector ${\bm{v}}\in{\mathbb{R}}^{n}$, the notation ${\bm{v}}_{-i}$ refers to the vector in ${\mathbb{R}}^{n-1}$ obtained by removing the $i^{\text{th}}$ entry of $\bm{v}$}. Likewise, ${\bm{X}}_{-i}$ denotes the matrix obtained by deleting the $i^{\text{th}}$ row of $\bm{X}$}.
Norms and asymptotic notation.
For a vector ${\bm{u}}\in{\mathbb{R}}^{n}$, $\left\|{\bm{u}}\right\|_{2}=\sqrt{{\bm{u}}^{\top}{\bm{u}}}$ denotes the Euclidean norm, and $\left\|{\bm{u}}\right\|_{\infty}=\max_{i\in[n]}|u_{i}|$ denotes the $\ell_{\infty}$ norm. The operator ${\rm tr}(\cdot)$ denotes the trace of a square matrix. For a matrix ${\bm{X}}\in{\mathbb{R}}^{n\times k}$, $\left\|{\bm{X}}\right\|_{F}=\sqrt{{\rm tr}({\bm{X}}^{\top}{\bm{X}})}$ denotes the Frobenius norm, and $\left\|{\bm{X}}\right\|_{2,\infty}=\max_{i\in[n]}\left\|{\bm{x}}_{i}\right\|_{2}$ denotes its $(2,\infty)$-operator norm. Let $\mathbb{N}$ denote the set of natural numbers. For any two functions $f,g\colon\mathbb{N}\to{\mathbb{R}}$, we write $f(n)=O(g(n))$ if there exist constants $C>0$ and $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$, $\lvert f(n)\rvert\leq C\lvert g(n)\rvert$.
Projection matrices.
Consider a full-rank matrix ${\bm{X}}\in{\mathbb{R}}^{n\times k}$ with $n\geq k$. We denote the projection matrix by $\boldsymbol{H}={\bm{X}}({\bm{X}}^{\top}{\bm{X}})^{-1}{\bm{X}}^{\top}$, and its diagonal entries by $h_{ii}={\bm{x}}_{i}^{\top}({\bm{X}}^{\top}{\bm{X}})^{-1}{\bm{x}}_{i}$, which represent the leverage scores of rows $i\in[n]$. The identity matrix is denoted by $\bm{I}$}. For $\lambda\geq 0$, the ridge projection matrix is $\boldsymbol{H}_{\lambda}={\bm{X}}({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{X}}^{\top}$, whose diagonal entries $h_{\lambda ii}={\bm{x}}_{i}^{\top}({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{x}}_{i}$ are the ridge leverage scores.
2 Finite Population Framework
We operate within the Neyman–Rubin potential outcomes framework (neyman1923application; rubin1974estimating) for a randomized controlled trial over a set of units indexed by $n]:=\{1,\ldots,n\}$. Each unit is assigned to either treatment ($T$) or control ($C$). The vector ${\bm{d}}\in{\mathbb{R}}^{n}$ denotes the treatment assignment, with $d_{i}=1$ if unit $i$ is in treatment and $d_{i}=0$ if it is in control. We also define ${\bm{z}}=2({\bm{d}}-1/2)$, so that $z_{i}=1$ for treated units and $z_{i}=-1$ otherwise. Finally, we define $\boldsymbol{q}$ as the vector with $i$-element equal to $q_{i}=p_{i}d_{i}+(1-p_{i})(1-d_{i})$.
For each unit $i$, the potential outcomes under treatment and control are denoted by $y^{(1)}_{i}$ and $y^{(0)}_{i}$, respectively. We treat these potential outcomes as fixed quantities; the only source of randomness in our setting is the treatment assignment. The observed outcome for unit $i$ is
$y_{i}=y^{(1)}_{i}d_{i}+y^{(0)}_{i}(1-d_{i}).$
We collect the observed and potential outcomes across all units in the vectors ${\bm{y}}=(y_{1},\ldots,y_{n})^{\top}$, ${\bm{y}}^{(1)}=(y^{(1)}_{1},\ldots,y^{(1)}_{n})^{\top}$, and ${\bm{y}}^{(0)}=(y^{(0)}_{1},\ldots,y^{(0)}_{n})^{\top}$.
The average treatment effect (ATE) is defined as
$\tau=\frac{1}{n}\sum_{i=1}^{n}\bigl(y^{(1)}_{i}-y^{(0)}_{i}\bigr)$.
Each unit $i\in[n]$ has a fixed vector of pretreatment covariates ${\bm{x}}_{i}\in{\mathbb{R}}^{k}$. The matrix ${\bm{X}}=({\bm{x}}_{1},\ldots,{\bm{x}}_{n})^{\top}\in{\mathbb{R}}^{n\times k}$ collects the covariates for all units. For simplicity of exposition, we assume throughout the paper that $\bm{X}$} has full column rank. However, since our estimators are based on ridge regression, the results extend directly to cases in which $\bm{X}$} is rank-deficient.
We consider two standard treatment assignment mechanisms. Under simple random assignment, each unit is independently assigned to treatment with probability $p_{i}\in(0,1)$. Let $\boldsymbol{p}=(p_{1},\ldots,p_{n})^{\top}$ denote the vector of assignment probabilities, and define
$m=\min_{i\in[n]}\min\{p_{i},\,1-p_{i}\}.$
Under complete random assignment, the number of treated units is fixed at $n_{T}\in[n-1]$, and the number of control units is $n_{C}=n-n_{T}$. Treatment is assigned by selecting uniformly at random from all possible subsets of $n_{T}$ units in the sample.
Under simple random assignment, the Horvitz–Thompson (HT) estimator (horvitz1952generalization) is defined as
$\widehat{\tau}_{\text{HT}}=\frac{1}{n}\sum_{i=1}^{n}\frac{d_{i}y_{i}}{p_{i}}\;-\;\frac{1}{n}\sum_{i=1}^{n}\frac{(1-d_{i})y_{i}}{1-p_{i}}.$
Under complete random assignment, the difference-in-means (DM) estimator is given by
$\widehat{\tau}_{\text{DM}}=\frac{1}{n_{T}}\sum_{i=1}^{n}d_{i}y_{i}\;-\;\frac{1}{n_{C}}\sum_{i=1}^{n}(1-d_{i})y_{i}.$
3 Regression Adjustment for Horvitz-Thompson Estimator
In this section, we first review some known results on the variance of the Horvitz–Thompson (HT) estimator. We then introduce LOORA-HT, our leave-one-out regression-adjusted version of the HT estimator (Section˜3.1). We establish its unbiasedness and derive an exact expression for its variance in the finite-population setting (Section˜3.2). Next, we analyze the asymptotic behavior of LOORA-HT (Section˜3.3) and prove its asymptotic efficiency (Section˜3.4). Finally, we describe our approach to variance estimation and confidence interval construction for LOORA-HT (Section˜3.5).
It is well known that, under simple random assignment, the classical Horvitz–Thompson (HT) estimator is unbiased for $\tau$ horvitz1952generalization. The variance of the HT estimator under simple random assignment is given by
$\frac{1}{n^{2}}\left\|{\boldsymbol{\mu}}\right\|_{2}^{2},$ (1)
where ${\boldsymbol{\mu}}=(\mu_{1},\ldots,\mu_{n})^{\top}$, and
$\mu_{i}=\sqrt{\frac{1-p_{i}}{p_{i}}}\,y^{(1)}_{i}+\sqrt{\frac{p_{i}}{1-p_{i}}}\,y^{(0)}_{i}$. (2)
Let $\breve{y}_{i}=y_{i}-{\bm{x}}_{i}^{\top}{\bm{b}}$ denote the covariate-adjusted outcomes, where $\bm{b}$} is a fixed vector of coefficients. The corresponding potential outcomes are
$\breve{y}^{(1)}_{i}=y^{(1)}_{i}-{\bm{x}}_{i}^{\top}{\bm{b}}\quad\text{and}\quad\breve{y}^{(0)}_{i}=y^{(0)}_{i}-{\bm{x}}_{i}^{\top}{\bm{b}}.$ (3)
Because $\breve{y}^{(1)}_{i}-\breve{y}^{(0)}_{i}=y^{(1)}_{i}-y^{(0)}_{i}$, it follows that the HT estimator applied to the adjusted outcomes $\breve{y}_{i}$ is also unbiased for $\tau$. Since $\bm{b}$} is fixed, applying (1) to the adjusted potential outcomes implies that the variance of the covariate-adjusted HT estimator is
$\frac{1}{n^{2}}\left\|{\boldsymbol{\mu}}-{\bm{R}}^{-1}{\bm{X}}{\bm{b}}\right\|_{2}^{2},$ (4)
where $\bm{R}$} is a diagonal matrix associated with the vector ${\bm{r}}\in{\mathbb{R}}^{n}$ defined by $r_{i}=\sqrt{p_{i}(1-p_{i})}$. Let $\widetilde{{\bm{X}}}={\bm{R}}^{-1}{\bm{X}}$. Then the variance in (4) is minimized by the choice of coefficients
$\boldsymbol{\beta}^{*}=\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in{\mathbb{R}}^{k}}\left\|{\boldsymbol{\mu}}-\widetilde{{\bm{X}}}{\bm{b}}\right\|_{2}^{2}.$ (5)
Equation (4) implies that regression adjustment can reduce the variance of the HT estimator. In practice, however, only one potential outcome is observed for each unit, so $\boldsymbol{\mu}$} cannot be computed directly, and the minimization problem in (5) cannot be solved exactly. We address this challenge in LOORA-HT by replacing $\boldsymbol{\mu}$} with a proxy vector equal to $\boldsymbol{\mu}$} in expectation.
3.1 Leave-One-Out Regression-Adjusted Horvitz-Thompson
Algorithm 1 LOORA-HT estimator for binary treatment experiments
1:Input: Covariates ${\bm{X}}\in{\mathbb{R}}^{n\times k}$, outcome vector ${\bm{y}}\in{\mathbb{R}}^{n}$, treatment assignment vector ${\bm{d}}\in\{0,1\}^{n}$, vector of treatment assignment probabilities $\boldsymbol{p}\in(0,1)^{n}$, regularization factor $\lambda\geq 0$.
2: Calculate ${\bm{r}}\in{\mathbb{R}}^{n}$ with $r_{i}=\sqrt{p_{i}(1-p_{i})}$ and $\widetilde{{\bm{X}}}={\bm{R}}^{-1}{\bm{X}}$.
3: Calculate $\boldsymbol{q}\in{\mathbb{R}}^{n}$ with
$q_{i}=\left\{\begin{array}[]{cl}p_{i}&\mbox{ if }d_{i}=1,\\ 1-p_{i}&\mbox{ if }d_{i}=0.\end{array}\right$.
4: Calculate ${\bm{z}}=2({\bm{d}}-1/2)\in{\mathbb{R}}^{n}$.
5: Calculate $\widetilde{{\bm{y}}}\in{\mathbb{R}}^{n}$ with
$\widetilde{y}_{i}=\frac{1}{q_{i}}\left(\frac{1-p_{i}}{p_{i}}\right)^{z_{i}/2}y_{i}.$
6: Set $S=0$.
7:for $i\in[n]$ do
8:   Set
$\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}=\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in\mathbb{R}^{k}}\left\|\widetilde{{\bm{y}}}_{-i}-\widetilde{{\bm{X}}}_{-i}{\bm{b}}\right\|_{2}^{2}+\lambda\left\|{\bm{b}}\right\|_{2}^{2}.$
9:   Set
$S=S+\frac{z_{i}}{q_{i}}(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}).$
10:end for
11:return
$\widehat{\tau}_{\textup{LHT}}=S/n.$
Algorithm˜1 describes LOORA-HT. Rather than using a single value for $\bm{b}$}, LOORA-HT applies a different vector of regression coefficients, $\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}$, for each unit $i$ in the sample. LOORA-HT constructs its adjustment vectors using a proxy vector $\widetilde{{\bm{y}}}$ that coincides with $\boldsymbol{\mu}$} in expectation,
$\widetilde{y}_{i}=\begin{cases}\dfrac{(1-p_{i})^{1/2}}{p_{i}^{3/2}}\,y_{i},&\text{if }d_{i}=1,\\[10.00002pt] \dfrac{p_{i}^{1/2}}{(1-p_{i})^{3/2}}\,y_{i},&\text{if }d_{i}=0.\end{cases}$
Each $\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}$ is obtained from a regression on the leave-one-out data, $(\widetilde{{\bm{y}}}_{-i},\widetilde{{\bm{X}}}_{-i})$. Leave-one-out coefficients can be particularly unstable for rows with high leverage scores (see Lemma˜B.1). To address this challenge, we use LOORA-HT ridge regression to mitigate the impact of high leverage scores. In Lemma˜2 below, we provide an upper bound on the ridge leverage scores as a function of the regularization parameter. For datasets with low leverage scores, the regularization parameter, $\lambda$, can be set to zero.
LOORA-HT is equal to
$\widehat{\tau}_{\textup{LHT}}=\frac{1}{n}\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}).$
3.2 The Variance of LOORA-HT
Let the ridge projection matrix of $\widetilde{{\bm{X}}}$ be
$\widetilde{\boldsymbol{H}}_{\lambda}=\widetilde{{\bm{X}}}(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}^{\top}$.
We denote its $(i,j)^{\text{th}}$ element by $\widetilde{h}_{\lambda ij}$ and its $i^{\text{th}}$ diagonal element by $\widetilde{h}_{\lambda ii}$. Next, we define the vector
${\bm{t}}=\frac{(1-\boldsymbol{p})^{2}{\bm{y}}^{(1)}-\boldsymbol{p}^{2}{\bm{y}}^{(0)}}{{\bm{r}}},$
which characterizes how far the vector $\widetilde{{\bm{y}}}$ deviates from $\boldsymbol{\mu}$}. More precisely,
$\widetilde{{\bm{y}}}-{\boldsymbol{\mu}}=\frac{{\bm{z}}{\bm{t}}}{\boldsymbol{q}},$
[…]
$\boldsymbol{\beta}_{\lambda}=\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in{\mathbb{R}}^{k}}\left\|{\boldsymbol{\mu}}-\widetilde{{\bm{X}}}{\bm{b}}\right\|_{2}^{2}+\lambda\left\|{\bm{b}}\right\|_{2}^{2}.$ (6)
The next theorem establishes the unbiasedness of LOORA-HT and provides an exact expression for its variance.
Theorem 1.
The LOORA-HT estimator (Algorithm˜1) is unbiased and its variance is given by the following formula:
$\frac{1}{n^{2}}\sum_{i=1}^{n}\frac{(\widetilde{{\bm{x}}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}-\mu_{i})^{2}}{(1-\widetilde{h}_{\lambda ii})^{2}}+\frac{1}{n^{2}}\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}(\widetilde{h}_{\lambda ij})^{2}\left(\frac{t_{j}}{r_{j}(1-\widetilde{h}_{\lambda ii})}+\frac{t_{i}}{r_{i}(1-\widetilde{h}_{\lambda jj})}\right)^{2}.$ (7)
Examining (7) closely, we observe that its first term mirrors the regression-adjusted version of the variance of the classical HT estimator in (4), and quantifies the variance reduction achieved by regression adjustment. The factor $(1-\widetilde{h}_{\lambda ii})^{-2}$ arises from the removal of a single row in the leave-one-out regressions. More specifically, by Lemma˜B.1, the first term of (7) can be written as
$\frac{1}{n^{2}}\sum_{i=1}^{n}(\widetilde{{\bm{x}}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}^{(-i)}-\mu_{i})^{2},$
where
$\boldsymbol{\beta}_{\lambda}^{(-i)}=\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in\mathbb{R}^{k}}\left\|{\boldsymbol{\mu}}_{-i}-\widetilde{{\bm{X}}}_{-i}{\bm{b}}\right\|_{2}^{2}+\lambda\left\|{\bm{b}}\right\|_{2}^{2}.$
Note that when a leverage score $\widetilde{h}_{\lambda ii}$ equals one, the variance becomes unbounded. In such high-leverage cases, it is essential to employ ridge regression with a sufficiently large regularization parameter $\lambda$. The following lemma provides an upper bound on the ridge leverage scores.
Lemma 2.
Let ${\bm{X}}\in{\mathbb{R}}^{n\times k}$, $c\geq 0$, and $\lambda=c\left\|{\bm{X}}\right\|_{2,\infty}^{2}$. Then, for all $i=1,\ldots,n$,
$h_{\lambda ii}\leq\frac{1}{1+c}.$
In our empirical analysis of within-subject experimental applications (Section˜6), we set $c=2$. It is well documented in the literature that units with high leverage scores can cause regression-adjusted estimators to yield unreliable estimates of the ATE (young2019channeling). Our LOORA estimators provide a systematic approach to mitigating this issue.
The second term of (7) arises from using the random vector $\widetilde{{\bm{y}}}$ in place of $\boldsymbol{\mu}$} in our regressions, since $\boldsymbol{\mu}$} is unobserved. As discussed earlier, ${\mathbb{E}}[\widetilde{{\bm{y}}}]={\boldsymbol{\mu}}$ and $\widetilde{{\bm{y}}}={\boldsymbol{\mu}}+\frac{{\bm{z}}{\bm{t}}}{\boldsymbol{q}}$. The second term of (7) thus captures the error introduced by estimating $\boldsymbol{\beta}_{\lambda}^{(-i)}$ with $\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}$, the solution to the regression based on $\widetilde{{\bm{y}}}$.
We next provide a loose upper bound for this term to show that it scales as $k/n^{2}$. Let
$\widetilde{{\bm{X}}}_{\lambda}=\begin{bmatrix}\widetilde{{\bm{X}}}\\[2.0pt] \sqrt{\lambda}\,{\bm{I}}\end{bmatrix}$.
It follows that
$\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}\widetilde{h}_{\lambda ij}^{2}$ $\leq\frac{1}{2}\left\|\widetilde{\boldsymbol{H}}_{\lambda}\right\|_{F}^{2}\leq\frac{1}{2}\left\|\widetilde{{\bm{X}}}_{\lambda}(\widetilde{{\bm{X}}}_{\lambda}^{\top}\widetilde{{\bm{X}}}_{\lambda})^{-1}\widetilde{{\bm{X}}}_{\lambda}^{\top}\right\|_{F}^{2}=\frac{k}{2},$
where the second inequality holds because $\widetilde{\boldsymbol{H}}_{\lambda}$ is an $(n\times n)$ block of the projection matrix associated with $\widetilde{{\bm{X}}}_{\lambda}$, and the final equality holds since the rank of $\widetilde{{\bm{X}}}_{\lambda}$ is $k$, and the trace of a projection matrix equals its rank.
Therefore, the second term of (7) is bounded by
$\frac{2k}{n^{2}}\left\|(1-\widetilde{\boldsymbol{h}}_{\lambda})^{-1}\right\|_{\infty}^{2}\left\|{\bm{t}}/{\bm{r}}\right\|_{\infty}^{2},$
where $\widetilde{\boldsymbol{h}}_{\lambda}$ denotes the vector of diagonal entries of $\widetilde{\boldsymbol{H}}_{\lambda}$. That is, for a fixed number of covariates, the second term of (7) scales as $1/n^{2}$ as $n$ increases, and is typically much smaller than the first term, which scales as $1/n$.
In particular, for $\lambda=\left\|{\bm{X}}\right\|_{2,\infty}^{2}$, by Lemma˜2, the variance of LOORA-HT is bounded by
$\frac{4}{n^{2}}\sum_{i=1}^{n}(\widetilde{{\bm{x}}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}-\mu_{i})^{2}+\frac{8k}{n^{2}}\left\|{\bm{t}}/{\bm{r}}\right\|_{\infty}^{2}\,.$
Increasing the value of $\lambda$ decreases the second term of (7). It also reduces the magnitude of $(1-\widetilde{h}_{\lambda ii})^{-2}$, but at the same time increases the quantity
$\frac{1}{n^{2}}\sum_{i=1}^{n}(\widetilde{{\bm{x}}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}-\mu_{i})^{2}.$
Since the vector $\mu$ is not observed, one cannot search for the optimal value of $\lambda$.
We next study the asymptotic behavior of the LOORA-HT estimator. Under mild regularity and uniform boundedness assumptions, the variance of the LOORA-HT estimator with $\lambda=0$ converges as follows:
${\rm Var}\bigl(\sqrt{n}\,(\widehat{\tau}_{\textup{LHT}}-\tau)\bigr)\;\longrightarrow\;\lim_{n\to\infty}\frac{1}{n}\left\|\widetilde{{\bm{X}}}\boldsymbol{\beta}^{*}-{\boldsymbol{\mu}}\right\|_{2}^{2},$
where $\boldsymbol{\beta}^{*}$ is the oracle coefficient vector defined in (5). This limiting variance coincides with the asymptotic variance of the classical regression adjustment estimator with interaction terms (lin2013agnostic).
3.3 Asymptotic Normality of LOORA-HT
In this section, we establish the asymptotic normality of the LOORA-HT estimator. We adopt the standard asymptotic framework under the Neyman–Rubin potential outcomes model, in which the sequence of units, together with their covariates and potential outcomes, is viewed as a deterministic infinite population. Formally, let $({\bm{x}}_{i},y_{i}^{(1)},y_{i}^{(0)})_{i\in\mathbb{N}}$ denote this sequence, and consider the truncated sample $({\bm{x}}_{i},y_{i}^{(1)},y_{i}^{(0)})_{i\in[n]}$ as $n\to\infty$. The estimator is analyzed on this growing sequence.
We assume that the number of covariates is fixed and impose a set of regularity conditions ensuring that covariates and potential outcomes are uniformly bounded and well-behaved in the limit. Relative to lin2013agnostic, we introduce two additional restrictions: (i) the covariates have uniformly bounded norms, and (ii) the individual treatment effects are uniformly bounded. The first condition guarantees that the ridge leverage scores converge uniformly to zero, while the second ensures that the additional variance component—namely, the second term of (7)—vanishes asymptotically. Note that these assumptions are mild, as otherwise the covariates and individual treatment effects would diverge as $n$ increases.
  1. Assumption 1.
    There exists a finite constant $L<\infty$ such that, for all $i\in\mathbb{N}$,
    $\left\|{\bm{x}}_{i}\right\|_{2}\leq L,\qquad\left|y_{i}^{(1)}-y_{i}^{(0)}\right|\leq L.$
    Moreover, for all $n\in\mathbb{N}$ and $j\in[k]$,
    $\frac{1}{n}\sum_{i=1}^{n}(y_{i}^{(1)})^{4}\leq L,\qquad\frac{1}{n}\sum_{i=1}^{n}(y_{i}^{(0)})^{4}\leq L,\qquad\frac{1}{n}\sum_{i=1}^{n}x_{ij}^{4}\leq L.$
  2. Assumption 2.
    Let $\bm{X}$} denote the $n\times k$ matrix whose $i$-th row is ${\bm{x}}_{i}$. Then $n^{-1}{\bm{X}}^{\top}{\bm{X}}$ converges to a bounded, positive definite matrix. Furthermore, the population means of $y_{i}^{(1)}$, $y_{i}^{(0)}$, $(y_{i}^{(1)})^{2}$, $(y_{i}^{(0)})^{2}$, $y_{i}^{(1)}y_{i}^{(0)}$, $y_{i}^{(1)}{\bm{x}}_{i}$, and $y_{i}^{(0)}{\bm{x}}_{i}$ converge to finite limits. For instance,
    $\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}y_{i}^{(1)}{\bm{x}}_{i}$
    exists and is finite.
  3. Assumption 3.
    For simple random assignment (used with LOORA-HT), let $p_{i}$ denote the treatment probability for unit $i$. We assume $m:=\inf_{i\in\mathbb{N}}\min\{p_{i},1-p_{i}\}$ is bounded away from zero. For complete random assignment (used with LOORA-DM), we assume that the treatment proportion $n_{T}/n$ converges to a limit $p_{T}$ satisfying $0<p_{T}<1$, which is bounded away from both 0 and $1$.
Recall that the outcome for unit $i$ is adjusted using the coefficient vector
$\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}=\arg\min_{{\bm{b}}\in\mathbb{R}^{d}}\left\|\widetilde{{\bm{X}}}_{-i}{\bm{b}}-\widetilde{{\bm{y}}}_{-i}\right\|_{2}^{2}+\lambda\left\|{\bm{b}}\right\|_{2}^{2}.$ (8)
Furthermore, define the population limit
$\boldsymbol{\beta}^{*}_{\lambda}:=\lim_{n\to\infty}\arg\min_{{\bm{b}}\in\mathbb{R}^{d}}\left\|\widetilde{{\bm{X}}}{\bm{b}}-{\boldsymbol{\mu}}\right\|_{2}^{2}+\lambda\left\|{\bm{b}}\right\|_{2}^{2}.$ (9)
The next lemma establishes the existence of $\boldsymbol{\beta}^{*}_{\lambda}$ and the convergence in probability of the leave-one-out estimators $\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$ to this limit.
Lemma 3.
Suppose the potential outcomes, covariates, and treatment assignment probabilities for the LOORA-HT estimator satisfy Assumptions˜1, 2 and 3. Let $\lambda\geq 0$ be fixed. Then the limit $\boldsymbol{\beta}^{*}_{\lambda}$ defined in (9) exists, and for all $i\in\mathbb{N}$,
$\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}\;\xrightarrow{p}\;\boldsymbol{\beta}^{*}_{\lambda}.$
The previous lemma establishes that the leave-one-out coefficients $\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$ converge in probability to their population limit $\boldsymbol{\beta}^{*}_{\lambda}$. This result ensures that, asymptotically, each unit’s adjusted outcome behaves as if it were constructed using the deterministic population parameter $\boldsymbol{\beta}^{*}_{\lambda}$ rather than the sample-dependent estimate $\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$. Consequently, the random variation introduced by finite-sample estimation of regression adjustments vanishes in large samples. Building on this observation, we can now characterize the asymptotic distribution of the LOORA-HT estimator itself.
Theorem 4.
Under Assumptions˜1, 2 and 3, $\sqrt{n}({\widehat{\tau}}_{\textup{LHT}}-\tau)$ converges in distribution to a Gaussian random variable with mean 0 and variance
$\lim_{n\to\infty}\frac{1}{n}\left\|\widetilde{{\bm{X}}}\boldsymbol{\beta}_{\lambda}^{*}-{\boldsymbol{\mu}}\right\|_{2}^{2}.$
Theorem 4 establishes the large-sample distribution of the LOORA-HT estimator and identifies its asymptotic variance in closed form. This result shows that, under mild regularity conditions, the regression adjustment embedded in LOORA-HT not only preserves consistency but also yields an asymptotically efficient estimator whose variance matches the limiting expression $\lim_{n\to\infty}n^{-1}\|\widetilde{{\bm{X}}}\boldsymbol{\beta}_{\lambda}^{*}-{\boldsymbol{\mu}}\|_{2}^{2}$. In the next subsection, we build on this characterization to analyze the asymptotic efficiency of LOORA-HT and show that its limiting variance coincides with that of the classic regression-adjusted estimator with interaction terms.
3.4 Asymptotic Efficiency of LOORA-HT
We now examine the asymptotic efficiency of the LOORA-HT estimator. Theorem 4 established that $\widehat{\tau}_{\textup{LHT}}$ is asymptotically normal with a well-defined limiting variance. In this subsection, we show that under simple random assignment, the asymptotic variance of LOORA-HT coincides with that of the regression-adjusted estimator with interaction terms proposed by lin2013agnostic for complete random assignment. This result implies that LOORA-HT achieves the same asymptotic efficiency as Lin’s estimator, despite being constructed under a different assignment mechanism and without explicitly introducing interaction terms between treatment indicators and covariates.
The distinction between simple and complete random assignment is largely asymptotically negligible, but it matters for finite samples and for the structure of the adjustment. Lin’s specification augments the regression model with treatment–covariate interaction terms, effectively doubling the number of covariate coefficients to be estimated. While this ensures consistency and efficiency in large samples, it can also increase the dimensionality of the regression problem and lead to instability when the number of covariates is large relative to the sample size. By contrast, LOORA-HT achieves the same asymptotic efficiency through a leave-one-out regularized adjustment that avoids interactions and remains computationally stable even with high-dimensional covariates.
lin2013agnostic establishes the following asymptotic variance for the regression-adjusted estimator with interaction terms, denoted by $\widehat{\tau}_{\text{interact}}$:
$\lim_{n\to\infty}{\rm Var}\!\big(\sqrt{n}(\widehat{\tau}_{\text{interact}}-\tau)\big)=\frac{1-p_{T}}{p_{T}}\lim_{n\to\infty}\sigma_{T}^{2}+\frac{p_{T}}{1-p_{T}}\lim_{n\to\infty}\sigma_{C}^{2}+2\lim_{n\to\infty}\sigma_{TC},$ (10)
where $p_{T}$ is the limiting treatment proportion and
$\sigma_{T}^{2}$ $:=\frac{1}{n}\left\|({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}_{T}^{*}-({\bm{y}}^{(1)}-\overline{{\bm{y}}}^{(1)})\right\|_{2}^{2},$
$\sigma_{C}^{2}$ $:=\frac{1}{n}\left\|({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}_{C}^{*}-({\bm{y}}^{(0)}-\overline{{\bm{y}}}^{(0)})\right\|_{2}^{2},$
$\sigma_{TC}$ $:=\frac{1}{n}\Big[\big(({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}_{T}^{*}-({\bm{y}}^{(1)}-\overline{{\bm{y}}}^{(1)})\big)^{\top}\big(({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}_{C}^{*}-({\bm{y}}^{(0)}-\overline{{\bm{y}}}^{(0)})\big)\Big],$
and
$\boldsymbol{\beta}_{T}^{*}$ $:=\lim_{n\to\infty}\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}}\left\|({\bm{X}}-\overline{{\bm{X}}}){\bm{b}}-({\bm{y}}^{(1)}-\overline{{\bm{y}}}^{(1)})\right\|_{2}^{2},$
$\boldsymbol{\beta}_{C}^{*}$ $:=\lim_{n\to\infty}\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}}\left\|({\bm{X}}-\overline{{\bm{X}}}){\bm{b}}-({\bm{y}}^{(0)}-\overline{{\bm{y}}}^{(0)})\right\|_{2}^{2}.$
We now express the right-hand side of (10) in a form that facilitates comparison with the asymptotic variance of LOORA-HT. Straightforward algebra yields
$\frac{1-p_{T}}{p_{T}}\sigma_{T}^{2}+\frac{p_{T}}{1-p_{T}}\sigma_{C}^{2}+2\sigma_{TC}$
$=\frac{\big\|(1-p_{T})\big(({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}_{T}^{*}-({\bm{y}}^{(1)}-\overline{{\bm{y}}}^{(1)})\big)+p_{T}\big(({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}_{C}^{*}-({\bm{y}}^{(0)}-\overline{{\bm{y}}}^{(0)})\big)\big\|_{2}^{2}}{n\,p_{T}(1-p_{T})}$
$=\frac{\big\|({\bm{X}}-\overline{{\bm{X}}})\!\big((1-p_{T})\boldsymbol{\beta}_{T}^{*}+p_{T}\boldsymbol{\beta}_{C}^{*}\big)-\big((1-p_{T})({\bm{y}}^{(1)}-\overline{{\bm{y}}}^{(1)})+p_{T}({\bm{y}}^{(0)}-\overline{{\bm{y}}}^{(0)})\big)\big\|_{2}^{2}}{n\,p_{T}(1-p_{T})}$
$=\frac{1}{n}\big\|({\bm{X}}-\overline{{\bm{X}}})\big(\sqrt{\tfrac{1-p_{T}}{p_{T}}}\boldsymbol{\beta}_{T}^{*}+\sqrt{\tfrac{p_{T}}{1-p_{T}}}\boldsymbol{\beta}_{C}^{*}\big)-({\boldsymbol{\mu}}-\overline{{\boldsymbol{\mu}}})\big\|_{2}^{2}.$
Moreover, one can verify that
$\sqrt{\tfrac{1-p_{T}}{p_{T}}}\boldsymbol{\beta}_{T}^{*}+\sqrt{\tfrac{p_{T}}{1-p_{T}}}\boldsymbol{\beta}_{C}^{*}=\lim_{n\to\infty}\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}}\left\|({\bm{X}}-\overline{{\bm{X}}}){\bm{b}}-({\boldsymbol{\mu}}-\overline{{\boldsymbol{\mu}}})\right\|_{2}^{2}.$
To connect this expression with LOORA-HT, consider the case where $p_{i}=p_{T}$ for all $i\in\mathbb{N}$. In this setting,
$\widetilde{{\bm{X}}}(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}})^{-1}\widetilde{{\bm{X}}}^{\top}{\boldsymbol{\mu}}={\bm{X}}({\bm{X}}^{\top}{\bm{X}})^{-1}{\bm{X}}^{\top}{\boldsymbol{\mu}}.$ (11)
To compare the asymptotic variance of LOORA-HT with that of $\widehat{\tau}_{\text{interact}}$, we set $\lambda=0$. Using standard calculations (see, e.g., Appendix A of ghadiri2023finite), if we apply Algorithm˜1 with the matrix $\big[\,\widetilde{{\bm{X}}}-\overline{\widetilde{{\bm{X}}}}\,\big|\,\textbf{1}\,\big],$ that is, centering the covariates and including an intercept, then the variance of $\sqrt{n}(\widehat{\tau}_{\textup{LHT}}-\tau)$ converges to
$\lim_{n\to\infty}\frac{1}{n}\min_{{\bm{b}}}\left\|(\widetilde{{\bm{X}}}-\overline{\widetilde{{\bm{X}}}}){\bm{b}}-({\boldsymbol{\mu}}-\overline{{\boldsymbol{\mu}}})\right\|_{2}^{2}=\lim_{n\to\infty}\frac{1}{n}\min_{{\bm{b}}}\left\|({\bm{X}}-\overline{{\bm{X}}}){\bm{b}}-({\boldsymbol{\mu}}-\overline{{\boldsymbol{\mu}}})\right\|_{2}^{2},$
where the equality follows from (11). Therefore, the asymptotic variance of LOORA-HT is equal to that of $\widehat{\tau}_{\text{interact}}$.
This equivalence demonstrates that LOORA-HT attains the same asymptotic efficiency as the regression-adjusted estimator with interaction terms of lin2013agnostic, even though it is constructed under simple random assignment and does not rely on explicitly modeling treatment–covariate interactions. In the next subsection, we develop valid and asymptotically correct confidence intervals for LOORA-HT.
3.5 Confidence Intervals for LOORA-HT
We now describe how to estimate the variance of the LOORA-HT estimator and construct corresponding confidence intervals. We treat $\widehat{\tau}_{\textup{LHT}}$ as a two-step estimator and show that the consistency of the second step does not depend on the consistency of the first-step parameter estimates. This implies that our variance estimator remains valid even when the first-step ridge coefficients $\widehat{\boldsymbol{\beta}}_{\lambda}$ are not estimated consistently. To estimate the variance in the second step, we employ the Huber–White (or “sandwich”) estimator of asymptotic variance.
Recall that LOORA-HT estimates the average treatment effect as
$\widehat{\tau}_{\textup{LHT}}=\frac{1}{n}\sum_{i=1}^{n}z_{i}\!\left(\frac{y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}}{q_{i}}\right)=\frac{1}{n}\sum_{i=1}^{n}z_{i}\!\left(\frac{y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}}{q_{i}(1-\widetilde{h}_{\lambda ii})}\right),$
where the second equality follows from Lemma˜B.1. Hence $\widehat{\tau}_{\textup{LHT}}$ can be interpreted as the coefficient from regressing the vector
$\left(\frac{y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}}{q_{i}(1-\widetilde{h}_{\lambda ii})}\right)_{i\in[n]}\quad\text{on}\quad{\bm{z}}=(z_{1},\ldots,z_{n})^{\top}.$
This representation makes it clear that $\widehat{\tau}_{\textup{LHT}}$ is a two-step estimator: in the first step, we obtain $\widehat{\boldsymbol{\beta}}_{\lambda}$ from a ridge regression of outcomes on covariates, and in the second step, we estimate the coefficient from a regression of the residualized outcomes on the treatment indicator. Although in practice we compute only the leave-one-out coefficients $\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}$, the residuals computed using $\widehat{\boldsymbol{\beta}}_{\lambda}$ and scaling with $(1-\widetilde{h}_{\lambda ii})^{-1}$ are equivalent by Lemma˜B.1.
Importantly, the consistency of $\widehat{\tau}_{\textup{LHT}}$ does not depend on whether $\widehat{\boldsymbol{\beta}}_{\lambda}$ consistently estimates the population limit $\boldsymbol{\beta}^{*}_{\lambda}$. Indeed, by construction,
${\mathbb{E}}\!\left[z_{i}\!\left(\frac{y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}}{q_{i}}\right)\right]=y_{i}^{(1)}-y_{i}^{(0)},$
so that $\widehat{\tau}_{\textup{LHT}}$ remains unbiased for $\tau$ and hence consistent under random assignment.
Huber–White variance and confidence intervals.
For this univariate regression, the residual for unit $i$ is
$\widehat{r}_{i}:=\frac{y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}}{q_{i}(1-\widetilde{h}_{\lambda ii})}-z_{i}\,\widehat{\tau}_{\textup{LHT}}.$
Let ${\bm{Z}}=(z_{1},\ldots,z_{n})^{\top}$ and $\widehat{{\bm{R}}}=(\widehat{r}_{1},\ldots,\widehat{r}_{n})^{\top}$. The Huber–White (HC0) variance estimator for the slope is
$\widehat{V}_{\textup{HW}}=({\bm{Z}}^{\top}{\bm{Z}})^{-1}{\bm{Z}}^{\top}\operatorname{diag}(\widehat{{\bm{R}}}\odot\widehat{{\bm{R}}}){\bm{Z}}({\bm{Z}}^{\top}{\bm{Z}})^{-1}=\frac{\sum_{i=1}^{n}z_{i}^{2}\,\widehat{R}_{i}^{2}}{(\sum_{i=1}^{n}z_{i}^{2})^{2}}.$
Since $z_{i}\in\{\pm 1\}$, this simplifies to $\widehat{V}_{\textup{HW}}=n^{-2}\sum_{i=1}^{n}\widehat{r}_{i}^{2}$. This estimator consistently estimates the asymptotic variance derived in Theorem 4 under mild moment conditions, and the “sandwich” form adjusts for the heteroskedasticity induced by unit-specific $p_{i}$ and the first-step estimation.
Finally, a $(1-\alpha)$ confidence interval for the average treatment effect is
$\widehat{\tau}_{\textup{LHT}}\;\pm\;z_{1-\alpha/2}\sqrt{\widehat{V}_{\textup{HW}}},$
where $z_{1-\alpha/2}$ is the $(1-\alpha/2)$ quantile of the standard normal distribution. This interval is asymptotically valid under the regularity conditions stated in Section 3.3.
4 Regression Adjustment for Difference-in-Means Estimator
We now consider the difference-in-means (DM) estimator under complete random assignment, where exactly $n_{T}$ units are assigned to treatment and $n_{C}=n-n_{T}$ to control. Throughout this subsection, the signal vector used in the Horvitz–Thompson analysis is redefined. Specifically,
$\widetilde{{\boldsymbol{\mu}}}\;:=\;n_{C}\,{\bm{y}}^{(1)}\;+\;n_{T}\,{\bm{y}}^{(0)},$
and all formulas below are written in terms of $\widetilde{{\boldsymbol{\mu}}}$ rather than the $\boldsymbol{\mu}$} used earlier.
The DM estimator
$\widehat{\tau}_{\textup{DM}}\;=\;\frac{1}{n_{T}}\sum_{i=1}^{n}d_{i}y_{i}\;-\;\frac{1}{n_{C}}\sum_{i=1}^{n}(1-d_{i})y_{i}$
is unbiased for the average treatment effect under complete random assignment. Its finite-population variance is given by
${\rm Var}(\widehat{\tau}_{\textup{DM}})\;=\;\frac{S_{n}^{2}(y^{(1)})}{n_{T}}\;+\;\frac{S_{n}^{2}(y^{(0)})}{n_{C}}\;-\;\frac{S_{n}^{2}(\tau)}{n},$
where $S_{n}^{2}(\cdot)$ denotes finite-population variances computed with the (n1)(n-1) divisor (see, e.g., imbens2015causal). Equivalently—and most convenient for our regression-adjustment analysis—this variance admits the representation
${\rm Var}\!\big(\widehat{\tau}_{\textup{DM}}\big)\;=\;\frac{1}{\,n_{T}\,n_{C}\,n\,(n-1)\,}\;\big\|\,\widetilde{{\boldsymbol{\mu}}}-\overline{\widetilde{{\boldsymbol{\mu}}}}\,\big\|_{2}^{2},$ (12)
where $\overline{\widetilde{{\boldsymbol{\mu}}}}:=(n^{-1}\sum_{i=1}^{n}\widetilde{\mu}_{i})\boldsymbol{1}$.
For covariate-adjusted outcomes $\breve{y}_{i}({\bm{b}})=y_{i}-{\bm{x}}_{i}^{\top}{\bm{b}}$, with fixed ${\bm{b}}\in\mathbb{R}^{k}$, the adjusted DM estimator
$\frac{1}{n_{T}}\sum_{i=1}^{n}d_{i}\,\breve{y}_{i}({\bm{b}})\;-\;\frac{1}{n_{C}}\sum_{i=1}^{n}(1-d_{i})\,\breve{y}_{i}({\bm{b}})$
remains unbiased for $\tau$, and its variance is
${\rm Var}\!\big(\widehat{\tau}_{\textup{DM}}({\bm{b}})\big)\;=\;\frac{1}{\,n_{T}\,n_{C}\,n\,(n-1)\,}\;\big\|\,\big(\widetilde{{\boldsymbol{\mu}}}-{\bm{X}}{\bm{b}}\big)-\overline{\big(\widetilde{{\boldsymbol{\mu}}}-{\bm{X}}{\bm{b}}\big)}\,\big\|_{2}^{2},$ (13)
where $\overline{(\widetilde{{\boldsymbol{\mu}}}-{\bm{X}}{\bm{b}})}:=(n^{-1}\sum_{i=1}^{n}(\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}{\bm{b}}))\boldsymbol{1}$. Therefore, the minimum attainable variance is
$\frac{1}{\,n_{T}\,n_{C}\,n\,(n-1)\,}\;\min_{{\bm{b}}\in\mathbb{R}^{k}}\big\|\,({\bm{X}}-\overline{{\bm{X}}})\,{\bm{b}}-(\widetilde{{\boldsymbol{\mu}}}-\overline{\widetilde{{\boldsymbol{\mu}}}})\,\big\|_{2}^{2}.$ (14)
However, since $\widetilde{{\boldsymbol{\mu}}}$ is not observable—each unit reveals only one potential outcome—this variance cannot be achieved in practice. In what follows, we construct an estimable surrogate with the correct conditional expectation and develop a leave-one-out regression-adjusted DM estimator (LOORA-DM), together with its finite-sample variance expression and asymptotic properties under complete random assignment.
4.1 Leave-One-Out Regression-Adjusted Difference-in-Means
Similar to LOORA-HT, our proposed leave-one-out regression-adjusted difference-in-means (LOORA-DM) estimator (Algorithm˜2) computes a distinct coefficient vector $\bm{b}$} for each unit $i$. A key difference, however, is that the regressed vector $\widetilde{{\bm{y}}}$ also differs across units. Under complete random assignment, the treatment assignments are not independent, so the regressed vector must be scaled differently depending on whether unit $i$ belongs to the treatment or control group in order to ensure that it has the correct expectation.
Algorithm 2 LOORA-DM estimator
1:Input: Covariates ${\bm{X}}\in{\mathbb{R}}^{n\times k}$, outcome vector ${\bm{y}}\in{\mathbb{R}}^{n}$, treatment assignment vector ${\bm{d}}\in\{0,1\}^{n}$ with $n_{T}=\sum_{i=1}^{n}d_{i}$ and $n_{C}=n-n_{T}$, regularization factor $\lambda\geq 0$.
2:Calculate $\bm{v}$} with $v_{i}=1/n_{T}$ if $d_{i}=1$, and $v_{i}=1/n_{C}$, otherwise.
3:Calculate ${\bm{f}}_{T}\in{\mathbb{R}}^{n}$ with $f_{Ti}=\frac{1}{n_{T}(n_{T}-1)}$ if $d_{i}=1$, and $f_{Ti}=\frac{1}{n^{2}_{C}}$ otherwise.
4:Calculate ${\bm{f}}_{C}\in{\mathbb{R}}^{n}$ with $f_{Ci}=\frac{1}{n_{T}^{2}}$ if $d_{i}=1$, and $f_{Ci}=\frac{1}{n_{C}(n_{C}-1)}$ otherwise.
5:Calculate ${\bm{z}}=2({\bm{d}}-1/2)\in\mathbb{R}^{n}$.
6:Set $S=0$.
7:for $i\in[n]$ do
8:  if $d_{i}=1$ then
9:   Let ${\bm{f}}^{(-i)}={\bm{f}}_{T}$.
10:  else
11:   Let ${\bm{f}}^{(-i)}={\bm{f}}_{C}$.
12:  end if
13:  Set
$\widetilde{{\bm{y}}}^{(-i)}=\frac{n_{T}n_{C}(n-1)}{n}{\bm{f}}^{(-i)}{\bm{y}}.$
14:  Set
$\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}=\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}}\left\|{\bm{X}}_{-i}{\bm{b}}-\widetilde{{\bm{y}}}^{(-i)}_{-i}\right\|_{2}^{2}+\lambda\left\|{\bm{b}}\right\|_{2}^{2}\,.$
15:  Set
$S=S+v_{i}z_{i}(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}^{(-i)}).$
16:end for
17:return
$\widehat{\tau}=S.$
More precisely, for each unit $i$, we define
$\widetilde{{\bm{y}}}^{(-i)}\;=\;\frac{n_{T}n_{C}(n-1)}{n}\,{\bm{f}}^{(-i)}{\bm{y}},$
where the components of ${\bm{f}}^{(-i)}$ are given by
$f^{(-i)}_{j}=\begin{cases}\dfrac{1}{n_{T}(n_{T}-1)},&\text{if }d_{j}=1,\\[10.00002pt] \dfrac{1}{n_{C}^{2}},&\text{if }d_{j}=0,\end{cases}\qquad\text{if }d_{i}=1,$
and
$f^{(-i)}_{j}=\begin{cases}\dfrac{1}{n_{T}^{2}},&\text{if }d_{j}=1,\\[10.00002pt] \dfrac{1}{n_{C}(n_{C}-1)},&\text{if }d_{j}=0,\end{cases}\qquad\text{if }d_{i}=0.$
The coefficient vector $\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$ is then obtained by regressing $\widetilde{{\bm{y}}}^{(-i)}_{-i}$ on ${\bm{X}}_{-i}$ using ridge regression. Finally, the outcome of unit $i$ is adjusted by ${\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$.
In the next subsection, we establish that LOORA-DM is unbiased for the average treatment effect and derive an exact expression for its variance.
4.2 The Variance of LOORA-DM
We now characterize the finite-sample variance of the leave-one-out regression-adjusted difference-in-means (LOORA-DM) estimator. The terms in the variance are more intricate here because, under complete random assignment, treatment assignments are not independent. Unbiasedness is also slightly more delicate to verify than in the Horvitz–Thompson setting, but it ultimately follows from the fact that $\widetilde{{\bm{y}}}^{(-i)}$ is constructed so that ${\mathbb{E}}\!\left[\widetilde{{\bm{y}}}^{(-i)}_{-i}\right]=\frac{\widetilde{{\boldsymbol{\mu}}}_{-i}}{n}$ for each $i$.
Theorem 5.
Under complete random assignment, the estimator obtained from Algorithm˜2 is unbiased and its variance is given by
$\frac{1}{n(n-1)n_{T}n_{C}}\sum_{i=1}^{n}\left(\frac{\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{1-h_{\lambda ii}}-\frac{1}{n}\sum_{j=1}^{n}\frac{\widetilde{\mu}_{j}-{\bm{x}}_{j}^{\top}\boldsymbol{\beta}_{\lambda}}{1-h_{\lambda jj}}\right)^{2}$
$\quad-\frac{2}{n^{2}(n-1)n_{T}n_{C}}\left(\sum_{\begin{subarray}{c}i,j\in[n]:\\ i\neq j\end{subarray}}\frac{(\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})\,h_{\lambda ij}\,\widetilde{\mu}_{i}}{(1-h_{\lambda ii})(1-h_{\lambda jj})}-\frac{1}{n-2}\sum_{\begin{subarray}{c}k\in[n]:\\ k\neq i,j\end{subarray}}\frac{(\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})\,h_{\lambda jk}\,\widetilde{\mu}_{k}}{(1-h_{\lambda ii})(1-h_{\lambda jj})}\right)$
$\quad+\begin{bmatrix}\widetilde{\bm{t}}^{(1)}\\ \widetilde{\bm{t}}^{(0)}\end{bmatrix}^{\!\top}\boldsymbol{Q}\begin{bmatrix}\widetilde{\bm{t}}^{(1)}\\ \widetilde{\bm{t}}^{(0)}\end{bmatrix}$,
where
$\widetilde{{\boldsymbol{\mu}}}$ $=n_{C}{\bm{y}}^{(1)}+n_{T}{\bm{y}}^{(0)},\qquad\boldsymbol{\beta}_{\lambda}=({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{X}}^{\top}\widetilde{{\boldsymbol{\mu}}}$,
$\widetilde{\bm{t}}^{(1)}$ $=n_{C}^{2}{\bm{y}}^{(1)}-n_{T}(n_{T}-1){\bm{y}}^{(0)},\qquad\widetilde{\bm{t}}^{(0)}=n_{C}(n_{C}-1){\bm{y}}^{(1)}-n_{T}^{2}{\bm{y}}^{(0)}.$
Entries of $\boldsymbol{Q}$ depend only on $n_{C}$, $n_{T}$, and the entries of the hat matrix $\boldsymbol{H}_{\lambda}$; its explicit form is provided in the appendix.
The first line in the variance expression can be viewed as the squared Euclidean norm of a vector after subtracting its sample mean. Specifically, it is the norm of
$\left(\frac{\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{1-h_{\lambda ii}}\right)_{i\in[n]}\;-\;\frac{1}{n}\sum_{j=1}^{n}\left(\frac{\widetilde{\mu}_{j}-{\bm{x}}_{j}^{\top}\boldsymbol{\beta}_{\lambda}}{1-h_{\lambda jj}}\right)\mathbf{1},$
scaled by the factor $n(n-1)n_{T}n_{C}]^{-1}$. The denominators $1-h_{\lambda ii}$ provide the leverage correction arising from leave-one-out adjustment. By contrast, in the Horvitz–Thompson case one must recenter the matrix and add an intercept explicitly to produce the analogous centering.
The second line collects cross-unit terms governed by the off-diagonal leverages $h_{\lambda ij}$. These cross-unit contributions did not appear in the Horvitz–Thompson analysis because units are independent under simple random assignment, whereas here the fixed treatment counts induce dependence across units under complete random assignment. The final quadratic form in $\boldsymbol{Q}$ bundles additional corrections that depend only on $(n_{T},n_{C})$ and the leverage structure encoded in $\boldsymbol{H}_{\lambda}$. For a fixed number of covariates, the first term scales as $O(1/n)$, while both the cross-unit term and the $\boldsymbol{Q}$-term scale as $O(1/n^{2})$; in particular, the second and third terms vanish asymptotically.
Taken together, Theorem 5 shows that the variance of LOORA-DM is dominated by the centered dispersion of the leverage-adjusted signal and that dependence induced by complete random assignment contributes only second-order terms. This decomposition will be useful when establishing asymptotic distributional results and efficiency comparisons later in the section.
4.3 Asymptotic Normality of LOORA-DM
This section is focused on the analysis of the large-sample behavior of the leave-one-out regression-adjusted difference-in-means (LOORA-DM) estimator under complete random assignment. Our objective is to establish that in large samples, the random leave-one-out coefficients used in LOORA-DM converge to a deterministic population limit and that the estimator itself satisfies a central limit theorem with a well-defined asymptotic variance.
Recall that for each unit $i$, the adjusted outcome is constructed using a coefficient vector estimated from all other units:
$\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$ $=\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in{\mathbb{R}}^{d}}\big\|{\bm{X}}_{-i}{\bm{b}}-\widetilde{{\bm{y}}}^{(-i)}_{-i}\big\|_{2}^{2}+\lambda\|{\bm{b}}\|_{2}^{2}.$ (15)
Here $\widetilde{{\bm{y}}}^{(-i)}$ rescales the observed outcomes in a manner that depends on whether the $i$th unit is assigned to treatment or control. This rescaling is necessary because, under complete random assignment, the treatment indicators are not independent across units, and the weights must therefore be chosen to preserve unbiasedness of the estimator.
To study the limiting behavior of these regression adjustments, define the population analogue
$\boldsymbol{\beta}^{*}_{\lambda}:=\lim_{n\to\infty}\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in{\mathbb{R}}^{d}}\left\|{\bm{X}}{\bm{b}}-\frac{\widetilde{{\boldsymbol{\mu}}}}{n}\right\|_{2}^{2}+\lambda\|{\bm{b}}\|_{2}^{2},$ (16)
where $\widetilde{{\boldsymbol{\mu}}}=n_{C}{\bm{y}}^{(1)}+n_{T}{\bm{y}}^{(0)}$. This limit represents the deterministic coefficient vector that best captures the relation between covariates and the signal component $\widetilde{{\boldsymbol{\mu}}}/n$ in the population. The next lemma shows that the leave-one-out coefficients converge in probability to this population limit.
Lemma 6.
Suppose the potential outcomes, covariates, and treatment assignment satisfy Assumptions˜1, 2 and 3, and let $\lambda\geq 0$ be fixed. Then $\boldsymbol{\beta}^{*}_{\lambda}$, as defined in (16), exists. Moreover, for all $i\in\mathbb{N}$,
$\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}\;\xrightarrow{p}\;\boldsymbol{\beta}^{*}_{\lambda}.$
This result implies that the random fluctuations in the estimated adjustment coefficients vanish asymptotically. Hence, in large samples, the behavior of LOORA-DM is equivalent to that of an “oracle” estimator that uses the deterministic coefficient $\boldsymbol{\beta}^{*}_{\lambda}$ for all units. This simplification enables a transparent analysis of the estimator’s asymptotic variance and distribution.
Substituting $\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$ with its probability limit $\boldsymbol{\beta}^{*}_{\lambda}$ yields the infeasible oracle estimator
$\widehat{\tau}^{*}_{\textup{LDM}}=\frac{1}{n_{T}}\sum_{i:d_{i}=1}\big(y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}^{*}_{\lambda}\big)-\frac{1}{n_{C}}\sum_{i:d_{i}=0}\big(y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}^{*}_{\lambda}\big).$
Because treatment is assigned without replacement under complete random assignment, the oracle estimator can be viewed as a contrast of two finite-population averages drawn by sampling without replacement. Applying Hájek’s finite-population central limit theorem (hajek1960limiting) yields the following result.
Theorem 7.
Under Assumptions˜1, 2 and 3, $\sqrt{n}\,(\widehat{\tau}_{\textup{LDM}}-\tau)$ converges in distribution to a Gaussian random variable with mean zero and variance
$\lim_{n\to\infty}\frac{1}{n}\big\|({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}_{\lambda}^{*}-({\boldsymbol{\mu}}-\overline{{\boldsymbol{\mu}}})\big\|_{2}^{2},$
where
${\boldsymbol{\mu}}=\sqrt{\frac{n_{C}}{n_{T}}}\,{\bm{y}}^{(1)}+\sqrt{\frac{n_{T}}{n_{C}}}\,{\bm{y}}^{(0)}.$
The variance expression above parallels that of the LOORA-HT estimator but differs in two key aspects. First, because the treatment indicators are dependent under complete random assignment, the centering of both $\bm{X}$} and $\boldsymbol{\mu}$} appears explicitly in the limiting variance. Second, the scaling of $\boldsymbol{\mu}$} reflects the finite-sample proportions of treated and control units, corresponding to the population fractions $n_{T}/n$ and $n_{C}/n$.
Together, Lemma 6 and Theorem 7 show that LOORA-DM achieves an asymptotically normal distribution with a variance determined by the projection of the population signal onto the null space of the covariate space. This characterization implies the asymptotic efficiency of LOORA-DM similar to LOORA-HT, see Section˜3.4.
4.4 Confidence Intervals for LOORA-DM
We now turn to inference for the LOORA-DM estimator. As in the case of LOORA-HT, we construct confidence intervals by viewing the estimator as the second coefficient in an auxiliary regression and estimating its variance using the heteroskedasticity-robust (Huber–White) variance formula. This approach treats LOORA-DM as a two-step estimator: the first step computes regression-adjusted residuals, and the second step averages these residuals across treatment and control groups.
Regression representation. Recall that the LOORA-DM estimator can be expressed as
$\widehat{\tau}_{\textup{LDM}}$ $=\frac{1}{n_{T}}\sum_{i:d_{i}=1}\big(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}\big)-\frac{1}{n_{C}}\sum_{i:d_{i}=0}\big(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}\big).$
Define the adjusted residual vector and design matrix
${\bm{u}}:=\left(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}\right)_{i=1}^{n},\qquad\boldsymbol{N}:=\begin{bmatrix}\boldsymbol{1}&~~{\bm{d}}\end{bmatrix}$,
where ${\bm{d}}=(d_{1},\ldots,d_{n})^{\top}$ is the treatment indicator vector. We now show that $\widehat{\tau}_{\textup{LDM}}$ equals the second coefficient from the least-squares projection of $\bm{u}$} onto $\boldsymbol{N}$.
Note that
$\boldsymbol{N}^{\top}\boldsymbol{N}=\begin{bmatrix}n&n_{T}\\ n_{T}&n_{T}\end{bmatrix},\qquad(\boldsymbol{N}^{\top}\boldsymbol{N})^{-1}=\frac{1}{n_{C}n_{T}}\begin{bmatrix}n_{T}&-n_{T}\\[1.99997pt] -n_{T}&n\end{bmatrix}$,
and
$\boldsymbol{N}^{\top}{\bm{u}}=\begin{bmatrix}\displaystyle\sum_{i=1}^{n}y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}\\[11.99998pt] \displaystyle\sum_{i:d_{i}=1}y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}\end{bmatrix}$.
Therefore,
$\big((\boldsymbol{N}^{\top}\boldsymbol{N})^{-1}\boldsymbol{N}^{\top}{\bm{u}}\big)_{2}$ $=-\frac{1}{n_{C}}\sum_{i=1}^{n}y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}+\frac{n}{n_{C}n_{T}}\sum_{i:d_{i}=1}y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}$
$=\frac{1}{n_{T}}\sum_{i:d_{i}=1}y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}-\frac{1}{n_{C}}\sum_{i:d_{i}=0}y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}.$
Thus $\widehat{\tau}_{\textup{LDM}}$ is precisely the second coefficient in the regression of $\bm{u}$} onto $\boldsymbol{N}$.
Huber–White variance estimator. Let
$\widehat{{\bm{V}}}_{\textup{HW}}=(\boldsymbol{N}^{\top}\boldsymbol{N})^{-1}\left(\boldsymbol{N}^{\top}\widehat{{\bm{S}}}\boldsymbol{N}\right)(\boldsymbol{N}^{\top}\boldsymbol{N})^{-1},\quad\text{where}\quad\widehat{{\bm{S}}}=\operatorname{diag}(\widehat{r}_{1}^{2},\ldots,\widehat{r}_{n}^{2}),$
and $\widehat{r}_{i}$ are the residuals from the regression,
$\widehat{r}_{i}=\frac{y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}}{1-\widetilde{h}_{\lambda ii}}-\widehat{\theta}-\widehat{\tau}_{\textup{LDM}}\,d_{i},$
with $(\widehat{\theta},\widehat{\tau}_{\textup{LDM}})$ the estimated coefficients from regressing $\bm{u}$} on $\boldsymbol{N}$. Then the estimated variance of $\widehat{\tau}_{\textup{LDM}}$ is given by the $(2,2)$ entry of $\widehat{{\bm{V}}}_{\textup{HW}}$:
$\widehat{{\rm Var}}(\widehat{\tau}_{\textup{LDM}})=e_{2}^{\top}\widehat{{\bm{V}}}_{\textup{HW}}e_{2},\qquad e_{2}=(0,1)^{\top}.$
This estimator coincides with the standard heteroskedasticity-robust (Huber–White) variance estimator for the slope coefficient in an OLS regression.
Inference and consistency. The regression formulation above implies that valid confidence intervals for the average treatment effect can be constructed using the estimated standard error
$\widehat{\sigma}=\sqrt{e_{2}^{\top}\widehat{{\bm{V}}}_{\textup{HW}}e_{2}},$
and normal critical values. Analogous to the case of LOORA-HT, the consistency of $\widehat{\tau}_{\textup{LDM}}$ does not depend on the consistency of $\widehat{\boldsymbol{\beta}}_{\lambda}$ itself: even if the first-stage ridge estimates are biased, the two-step estimator remains consistent for the average treatment effect under complete random assignment.
5 Discussion
The estimators developed in this paper contribute to a long line of work on regression adjustment and design-based inference in randomized experiments. A central goal of this literature is to reconcile two desiderata that are often in tension: finite-sample unbiasedness under the experimental design, and efficiency gains through regression-based covariate adjustment. Classical estimators such as the Horvitz–Thompson and difference-in-means achieve the former but not the latter, while regression-adjusted estimators improve efficiency at the cost of bias or instability in small samples. The LOORA estimators bridge this gap by providing finite-sample unbiased and asymptotically efficient estimators that remain robust to high-leverage covariates.
To highlight how the LOORA estimators relate to existing methods, Table 1 summarizes the key properties of several classical estimators under simple and complete randomization. The comparison focuses on unbiasedness, asymptotic efficiency relative to the benchmark of lin2013agnostic, and robustness to leverage. The table emphasizes that while classical regression adjustments can reduce variance, none of them simultaneously ensure all three properties.
Classical estimators. Under simple randomization, the Horvitz–Thompson (HT) estimator is exactly unbiased but can exhibit high variance. Regression adjustment aims to improve efficiency by controlling for observed covariates. The most commonly used adjusted estimator, $\widehat{\tau}_{\text{ADJ}}$, regresses $\bm{y}$} on $\bm{X}$} and the treatment indicator $\bm{d}$} with intercept, producing a consistent estimator only when treatment effects are homogeneous. The interacted version, $\widehat{\tau}_{\text{INT}}$, adds treatment-by-covariate interactions and is asymptotically efficient under both simple and complete randomization (lin2013agnostic). However, both estimators can become unstable when the covariate matrix contains influential or high-leverage observations, and neither is finitely unbiased under the randomization design (young2019channeling).
Leave-one-out regression adjustment. The LOORA framework modifies regression adjustment to preserve finite-sample unbiasedness without sacrificing efficiency. By using leave-one-out fitted values, the LOORA-HT and LOORA-DM estimators remove the bias introduced by in-sample residual reuse while maintaining asymptotic optimality. Under simple randomization, LOORA-HT achieves the same variance bound as the lin2013agnostic estimator but is exactly unbiased and robust to leverage. Under complete randomization, LOORA-DM extends these results to dependent assignment structures by appropriately rescaling the leave-one-out predictions. Both estimators therefore combine the finite-sample exactness of HT and DM with the asymptotic efficiency of regression adjustment.
Relation to recent work. Recent contributions have revisited the foundations of design-based inference from distinct perspectives. ghadiri2023finite develop a finite-sample unbiased regression-adjusted estimator under simple random assignment, using a cross-fitted formulation that motivates the present work. However, their approach does not attain the asymptotically efficient variance bound. harshaw2022design, in contrast, take a design-based perspective: rather than modifying the estimator, they optimize the randomization mechanism itself to minimize the variance of the unadjusted Horvitz–Thompson estimator. Finally, spiess2025optimal characterize the necessary conditions under which design-based estimators such as Horvitz–Thompson or difference-in-means are unbiased. For the HT estimator, these conditions require a leave-one-out structure, which our LOORA-HT estimator satisfies by construction. For the DM estimator, unbiasedness requires a leave-two-out structure; we show that LOORA-DM can be equivalently viewed as a leave-two-out estimator, thereby satisfying Spiess’s condition. Hence, our estimators are not only consistent with the theoretical framework of spiess2025optimal, but also provide constructive, closed-form realizations of unbiased estimators that satisfy these necessary design-based criteria. Detailed comparisons with these works are presented in the subsections that follow.
Summary of properties. As shown in Table 1, the LOORA-HT and LOORA-DM estimators uniquely satisfy all three desirable properties: they are finitely unbiased, asymptotically efficient, and robust to high-leverage covariate realizations under their respective randomization schemes. Classical regression-adjusted estimators achieve at most two of these properties. By contrast, the LOORA estimators combine the robustness of design-based estimators with the efficiency of regression adjustment in a single, tractable framework.
Table 1: Properties of estimators
unbiased attains lin2013agnostic robust to
asymptotic high
variance leverage
Under simple randomization:
$\widehat{\tau}_{\text{HT}}$
$\widehat{\tau}_{\text{LOORA-HT}}$
Under complete randomization:
$\widehat{\tau}_{\text{DM}}$
$\widehat{\tau}_{\text{ADJ}}$ (Classic Regression Adjustment)
$\widehat{\tau}_{\text{INT}}$ (Interacted Regression Adjustment lin2013agnostic)
$\widehat{\tau}_{\text{LOORA-DM}}$
5.1 Comparison with Cross-Fitted Regression-Adjusted HT
ghadiri2023finite propose a finite-sample unbiased regression-adjusted Horvitz–Thompson estimator based on a cross-fitting approach; that is, the sample is split into two groups, a regression vector is learned separately within each group, and the fitted vector from one group is used to adjust the outcomes of the other. They consider only the setting where $p_{i}=0.5$ for all $i$, and provide the following upper bound on the variance of their estimator:
$\frac{8(1+\varepsilon)}{n^{2}}\min_{{\bm{b}}\in\mathbb{R}^{d}}\Big(\left\|{\bm{X}}{\bm{b}}-{\boldsymbol{\mu}}\right\|_{2}^{2}+100\log(n/\delta)\,\zeta_{{\bm{X}}}^{2}\left\|{\bm{b}}\right\|_{2}^{2}\Big)+\frac{32d}{n^{2}}\left\|{\bm{y}}^{(1)}-{\bm{y}}^{(0)}\right\|_{\infty}^{2}.$ (17)
By Lemma˜2, with their choice of regularization parameter $\lambda=100\log(n/\delta)\,\zeta_{{\bm{X}}}^{2}$, the leverage scores satisfy $h_{i}({\bm{X}},\lambda)\leq\frac{1}{1+100\log(n/\delta)}$. Moreover, since all $p_{i}=0.5$, it follows that
$\max_{i\in[n]}\left\|\widetilde{{\bm{x}}}_{i}\right\|_{2}=2\max_{i\in[n]}\left\|{\bm{x}}_{i}\right\|_{2}.$
Hence, noting that $4\lambda=100\log(n/\delta)\,\zeta_{\widetilde{{\bm{X}}}}^{2}$, we have $h_{i}({\bm{X}},\lambda)=h_{i}(\widetilde{{\bm{X}}},4\lambda)$.
Now consider
$\boldsymbol{\beta}$ $=\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in\mathbb{R}^{d}}\Big(\left\|{\bm{X}}{\bm{b}}-{\boldsymbol{\mu}}\right\|_{2}^{2}+100\log(n/\delta)\,\zeta_{{\bm{X}}}^{2}\left\|{\bm{b}}\right\|_{2}^{2}\Big)$
$=\frac{1}{2}\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in\mathbb{R}^{d}}\Big(\left\|\widetilde{{\bm{X}}}{\bm{b}}-{\boldsymbol{\mu}}\right\|_{2}^{2}+100\log(n/\delta)\,\zeta_{\widetilde{{\bm{X}}}}^{2}\left\|{\bm{b}}\right\|_{2}^{2}\Big).$
The variance of our estimator, from Theorem˜1, is bounded by
$\frac{1}{n^{2}(1-\frac{1}{101})^{2}}\sum_{i=1}^{n}(\widetilde{{\bm{x}}}_{i}^{\top}(2\boldsymbol{\beta})-\mu_{i})^{2}+\frac{8d}{n^{2}}\left\|{\bm{t}}/{\bm{r}}\right\|_{\infty}^{2}$
$=\frac{1.0201}{n^{2}}\sum_{i=1}^{n}({\bm{x}}_{i}^{\top}\boldsymbol{\beta}-\mu_{i})^{2}+\frac{8d}{n^{2}}\left\|{\bm{y}}^{(1)}-{\bm{y}}^{(0)}\right\|_{\infty}^{2}.$
Relative to the bound of ghadiri2023finite in (17), the bound above improves the first term by approximately a factor of eight. It also eliminates the additional ridge regularization penalty term appearing in (17), and improves the final term by a factor of four.
Finally, it is worth noting that ghadiri2023finite do not provide an inference procedure or a method for constructing confidence intervals, whereas we provide a two-step estimator perspective that allows consistent variance estimation and valid confidence interval construction.
5.2 Comparison with Gram–Schmidt Random Walk Design
harshaw2022design propose an alternative, design-based approach to variance reduction. Rather than modifying the estimator, they optimize the randomization mechanism itself through the Gram–Schmidt walk (GSW) design, which ensures covariate balance while maintaining marginal assignment probabilities. While such algorithmic designs can achieve improved covariate balance in finite samples, they require using the covariates prior to the experiment when constructing the design. In contrast, under simple or complete random assignment, one can freely add or remove covariates, analyze different subgroups of the population, or draw additional samples even after the experiment has been conducted. This flexibility—together with the transparency and ease of implementation of standard randomization— makes simple and complete random designs more attractive in most empirical settings.
harshaw2022design show that their GSW design achieves the following variance bound for any choice of $\phi\in(0,1)$:
$\frac{1}{n^{2}}\min_{\boldsymbol{\beta}\in\mathbb{R}^{d}}\Big[\frac{1}{\phi}\left\|{\boldsymbol{\mu}}-{\bm{X}}\boldsymbol{\beta}\right\|_{2}^{2}+\frac{\zeta_{{\bm{X}}}^{2}}{1-\phi}\left\|\boldsymbol{\beta}\right\|_{2}^{2}\Big].$ (18)
We now compare this bound with that of our LOORA-HT estimator.
The second term in (18) and the corresponding regularization term in (7) are not directly comparable in general, although both scale on the order of $k/n^{2}$. Setting $\lambda=\zeta_{{\bm{X}}}^{2}/(1-\phi)$, Lemma˜2 implies that
$h_{i}(\widetilde{{\bm{X}}},\tfrac{\zeta_{\widetilde{{\bm{X}}}}^{2}}{1-\phi})\;\leq\;\frac{1-\phi}{2-\phi}.$
Consequently,
$\frac{1}{1-\big(h_{i}(\widetilde{{\bm{X}}},\tfrac{\zeta_{\widetilde{{\bm{X}}}}^{2}}{1-\phi})\big)^{2}}\;\leq\;(2-\phi)^{2}.$
For $\phi\in[0,(3-\sqrt{5})/2)$, we have $(2-\phi)^{2}\leq 1/\phi$. Hence, the scaling factor that appears in our variance bound is uniformly smaller than that of the GSW design. Moreover, the first term in (18) diverges as $\phi\to 0$, whereas the corresponding constant in the variance of LOORA-HT remains bounded by $4$.
Finally, when the leverage scores of $\widetilde{{\bm{X}}}$ are small—that is, $h_{i}(\widetilde{{\bm{X}}},0)$ is small for all ii—we may safely take $\lambda=0$ in our estimator. In contrast, for the GSW design, the smallest feasible regularization parameter is $\lambda=\zeta_{{\bm{X}}}^{2}$, so the variance bound cannot achieve the same limiting rate. This highlights a key practical advantage of LOORA-HT under standard randomization schemes: it achieves comparable or tighter variance guarantees without altering the experimental design itself.
5.3 Connections of LOORA-DM to Leave-Two-Out Estimators
spiess2025optimal provides a general characterization of unbiased design-based estimators for average treatment effects under arbitrary randomization schemes. He shows that for Horvitz–Thompson (HT) and difference-in-means (DM) type estimators, unbiasedness imposes specific structural constraints on how each unit’s outcome can depend on treatment assignments of other units. In particular, the HT estimator must have a leave-one-out structure—each adjusted outcome for unit $i$ may depend on the treatment assignments of all other units except ii—whereas the DM estimator must satisfy a stronger leave-two-out condition, meaning that each pairwise contribution between units $i$ and $j$ can depend only on the treatment assignments of units other than $i$ and $j$.
In what follows, we show that the LOORA-DM estimator satisfies exactly this leave-two-out property, and thus conforms to the necessary unbiasedness conditions established by spiess2025optimal.
We show that LOORA-DM estimator can be written as
$\widehat{\tau}_{\textup{LDM}}=\frac{1}{n_{T}n_{C}}\sum_{i<j}(d_{i}-d_{j})\big(y_{i}-y_{j}-\phi_{ij}({\bm{z}}_{-ij})\big),$
where ${\bm{z}}_{-ij}$ denotes the treatment assignments of all units except $i$ and $j$. Note that
$\frac{1}{n_{T}n_{C}}\sum_{i<j}(d_{i}-d_{j})(y_{i}-y_{j})=\frac{1}{n_{T}}\sum_{d_{i}=1}y_{i}-\frac{1}{n_{C}}\sum_{d_{i}=0}y_{i},$
so it suffices to characterize the adjustment term $\phi_{ij}({\bm{z}}_{-ij})$.
Adjustment for treated and control units.
For a treated unit $e$ (i.e., $d_{e}=1$), the adjustment in the LOORA-DM estimator is
$-\frac{1}{n_{T}}{\bm{x}}_{e}^{\top}({\bm{X}}_{-e}^{\top}{\bm{X}}_{-e}+\lambda{\bm{I}})^{-1}{\bm{X}}_{-e}^{\top}\widetilde{{\bm{y}}}^{(T)}_{-e}$, (19)
where
$\widetilde{y}^{(T)}_{\ell}=\begin{cases}\frac{n_{C}(n-1)}{(n_{T}-1)n}y^{(1)}_{\ell},&d_{\ell}=1,\\[2.58334pt] \frac{n_{T}(n-1)}{n_{C}n}y^{(0)}_{\ell},&d_{\ell}=0.\end{cases}$
For a control unit $e$ ($d_{e}=0$), the corresponding adjustment is
$+\frac{1}{n_{C}}{\bm{x}}_{e}^{\top}({\bm{X}}_{-e}^{\top}{\bm{X}}_{-e}+\lambda{\bm{I}})^{-1}{\bm{X}}_{-e}^{\top}\widetilde{{\bm{y}}}^{(C)}_{-e}$, (20)
with
$\widetilde{y}^{(C)}_{\ell}=\begin{cases}\frac{n_{C}(n-1)}{n_{T}n}y^{(1)}_{\ell},&d_{\ell}=1,\\[2.58334pt] \frac{n_{T}(n-1)}{(n_{C}-1)n}y^{(0)}_{\ell},&d_{\ell}=0.\end{cases}$
Pairwise representation.
Define a unified adjustment quantity
$\widetilde{y}_{\ell}=\begin{cases}\frac{n_{C}(n-1)}{(n_{T}-1)n}y^{(1)}_{\ell},&d_{\ell}=1,\\[2.58334pt] \frac{n_{T}(n-1)}{(n_{C}-1)n}y^{(0)}_{\ell},&d_{\ell}=0,\end{cases}$
and set
$\phi_{ij}({\bm{z}}_{-ij})={\bm{x}}_{i}^{\top}({\bm{X}}_{-i}^{\top}{\bm{X}}_{-i}+\lambda{\bm{I}})^{-1}{\bm{X}}_{-ij}^{\top}\widetilde{{\bm{y}}}_{-ij}-{\bm{x}}_{j}^{\top}({\bm{X}}_{-j}^{\top}{\bm{X}}_{-j}+\lambda{\bm{I}})^{-1}{\bm{X}}_{-ij}^{\top}\widetilde{{\bm{y}}}_{-ij}.$
This form makes explicit that $\phi_{ij}$ depends only on ${\bm{z}}_{-ij}$—that is, on the treatment assignments of all units other than $i$ and jj—and hence defines a leave-two-out structure.
Substituting this definition, we obtain
$-\frac{1}{n_{T}n_{C}}\sum_{i<j}(d_{i}-d_{j})\phi_{ij}({\bm{z}}_{-ij})={}$ $-\frac{1}{n_{T}n_{C}}\sum_{d_{i}=1}{\bm{x}}_{i}^{\top}({\bm{X}}_{-i}^{\top}{\bm{X}}_{-i}+\lambda{\bm{I}})^{-1}\sum_{d_{j}=0}{\bm{X}}_{-ij}^{\top}\widetilde{{\bm{y}}}_{-ij}$
$+\frac{1}{n_{T}n_{C}}\sum_{d_{i}=0}{\bm{x}}_{i}^{\top}({\bm{X}}_{-i}^{\top}{\bm{X}}_{-i}+\lambda{\bm{I}})^{-1}\sum_{d_{j}=1}{\bm{X}}_{-ij}^{\top}\widetilde{{\bm{y}}}_{-ij}.$
Simple algebra then shows that, for treated units ($d_{i}=1$), the right-hand side reduces exactly to (19), and for control units ($d_{i}=0$) to (20). Thus, $\widehat{\tau}_{\textup{LDM}}$ can indeed be expressed in the leave-two-out form above.
Interpretation.
This representation shows that LOORA-DM satisfies the structural unbiasedness conditions identified by spiess2025optimal. Specifically, the adjustment $\phi_{ij}({\bm{z}}_{-ij})$ for each pair $(i,j)$ depends only on the treatment assignments of the remaining units, ensuring that the contribution of each pair is conditionally unbiased given the rest of the assignment vector. Hence, while the LOORA-HT estimator satisfies the leave-one-out property required for Horvitz–Thompson estimators, LOORA-DM naturally satisfies the stronger leave-two-out property required for difference-in-means estimators. Our construction therefore provides a concrete, closed-form realization of the abstract unbiasedness conditions derived in spiess2025optimal.
6 Experiments
This section provides empirical evidence on the finite-sample performance of our proposed estimators. We use experimental data from two within-subject studies, which allow us to observe both treatment and control outcomes for each unit. These data make it possible to generate synthetic randomized experiments under controlled conditions, enabling a systematic comparison of estimator bias, efficiency, and confidence-interval coverage. Our goal is not to assess the substantive findings of the original studies, but rather to evaluate the statistical behavior of LOORA–HT and LOORA–DM relative to standard regression-adjusted estimators.
The datasets are drawn from allcott2015evaluating and mcdonald2025evaluating. A brief description of each is provided below, and our replication code—available online—contains full details for data processing and experimental design.
  1. 1.
    Statehood dataset. mcdonald2025evaluating study the persuasiveness of policy arguments using a repeated-measures, within-subject design. Respondents first record their opinions on a policy issue, then read one or more arguments either supporting or opposing the policy, and finally re-state their opinions after exposure to each argument. We use the portion of their data concerning opinions about granting statehood to the District of Columbia (DC). The outcome under control is the respondent’s opinion before reading any arguments; the outcome under treatment is the opinion after reading (i) an argument against DC statehood emphasizing political corruption, and (ii) an argument in favor highlighting taxation without representation. We restrict attention to male respondents aged 40–49. Covariates include one-hot encodings of six categorical variables: party affiliation, race, voter registration, political attentiveness, education, and ideology. The resulting dataset contains $36$ units and $32$ covariates.
  2. 2.
    Lightbulb dataset. allcott2015evaluating conduct an experiment on consumer preferences between two types of lightbulbs offered at varying relative prices. Respondents first choose between the two options, after which they receive information about the energy costs of each type and make the choice again. The initial choice forms the control outcome; the post-information choice forms the treatment outcome. Covariates are obtained by one-hot encoding categorical features from the original data, including employment status, renter status, U.S. region, metropolitan statistical area status, marital status, income, housing type, gender, and ethnicity. The final dataset contains $123$ units and $53$ covariates.
Dataset Method Bias STD RMSE CI coverage (95%) CI average length (95%)
Statehood REG 0.0504 0.5191 0.5216 0.7182 0.5783
Ridge-REG 0.0887 0.2708 0.2849 0.8956 0.4467
LOORA-HT 0.0010 0.3125 0.3125 0.9425 0.6274
LOORA-DM 0.0510 0.2823 0.2869 0.9461 0.5880
Lightbulb REG 0.0021 0.1054 0.1054 0.9115 0.1797
Ridge-Reg 0.0274 0.0855 0.0898 0.9198 0.1542
LOORA-HT 0.0000 0.0849 0.0849 0.9583 0.1746
LOORA-DM 0.0021 0.0799 0.0799 0.9619 0.1668
Table 2: Performance metrics for four methods on two datasets with treatment assignment probabilities correlated with covariates.
Because both datasets arise from within-subject designs, we observe for each individual both potential outcomes. This feature allows us to simulate many synthetic randomized experiments by masking one outcome per unit according to a prescribed randomization rule. We can then evaluate the finite-sample properties of each estimator—bias, standard deviation, root mean squared error (RMSE), and the average length and coverage of confidence intervals—under realistic covariate distributions.
We consider two treatment-assignment mechanisms. In both cases, assignments are independent across units. In the first design, the treatment probability varies with the covariates. Specifically, we draw a random Gaussian vector, compute for each unit $i$ the cosine similarity $c_{i}$ between this vector and its covariate vector, and set
$p_{i}=\max\!\left\{\min\!\left\{\frac{1+c_{i}}{2},\,0.8\right\},\,0.2\right\}.$
In the second design, the treatment probability is fixed at $p_{i}=0.5$ for all units.
Each synthetic experiment is repeated 100,000100{,}000 times. Tables 2 and 3 summarize the results. We compare our estimators, LOORA–HT and LOORA–DM, with two benchmarks: (i) the classical regression-adjusted estimator (REG) of lin2013agnostic, and (ii) a ridge-regularized variant (Ridge–REG).
Our estimators consistently achieve lower RMSE than REG, reflecting both lower variance and negligible bias. While Ridge–REG attains a similar RMSE, this occurs at the cost of a larger bias—often exceeding that of REG—because the regularization distorts finite-sample unbiasedness. Moreover, the confidence intervals derived from our estimators, using the Huber–White variance estimator discussed in Sections˜3.5 and 4.4, exhibit coverage rates remarkably close to the nominal level, whereas both REG and Ridge–REG tend to undercover substantially. We note that the bias of LOORA–DM in these experiments is somewhat larger than that of LOORA–HT, because the treatment assignments are generated under simple random assignment rather than complete random assignment. These findings confirm that leave-one-out regularization provides a principled way to stabilize regression adjustment without sacrificing unbiasedness or inferential validity.
7 Conclusion
This paper develops leave-one-out regression-adjusted estimators for treatment effects under simple and complete random assignment. We show that these estimators are exactly unbiased in finite populations, achieve the asymptotic efficiency of the interacted regression of lin2013agnostic, and admit closed-form expressions for their variance. Regularization through leave-one-out ridge adjustment ensures robustness to high-leverage observations and stabilizes inference even in small samples. Our theoretical results establish a design-based foundation for regression adjustment that unifies classical unbiasedness, modern efficiency results, and practical inference procedures. Empirical evaluations confirm the estimators’ excellent finite-sample performance across different designs.
Beyond the settings considered here, future work could extend our framework to stratified or covariate-adaptive designs, to clustered or networked experiments, and to high-dimensional or nonparametric adjustments. Such extensions would further connect design-based causal inference with modern statistical learning while preserving the transparency and interpretability that motivate experimental design in the first place.
Dataset Method Bias STD RMSE CI coverage (95%) CI average length (95%)
Statehood REG 0.0432 0.5100 0.5118 0.7259 0.5785
Ridge-REG 0.0559 0.2692 0.2749 0.8989 0.4515
LOORA-HT 0.0009 0.2856 0.2856 0.9531 0.5901
LOORA-DM 0.0017 0.2781 0.2781 0.9557 0.5849
Lightbulb REG 0.0023 0.1048 0.1048 0.9094 0.1783
Ridge-Reg 0.0297 0.0861 0.0911 0.9190 0.1533
LOORA-HT 0.0000 0.0822 0.0822 0.9616 0.1705
LOORA-DM 0.0000 0.0797 0.0797 0.9623 0.1671
Table 3: Performance metrics for four methods on two datasets with treatment assignment probabilities all equal to $0.5$.
Appendix A Linear Algebra Tools
Lemma A.1 (alaoui2015fast; fahrbach2022subquadratic).
Let ${\bm{U}}{\boldsymbol{\Sigma}}{\bm{V}}$ be the compact SVD of $\bm{X}$}. Let $r$ be the rank of $\bm{X}$} and $\sigma_{1},\ldots,\sigma_{r}$ be its singular values. Then
$h_{\lambda ii}=\sum_{j=1}^{r}\frac{\sigma_{j}^{2}u_{ij}^{2}}{\sigma_{j}^{2}+\lambda}\,.$
Proof of Lemma˜2.
Let ${\bm{U}}{\boldsymbol{\Sigma}}{\bm{V}}$ be the compact SVD of $\bm{X}$} and $r$ be its rank. Then by Lemma˜A.1,
$h_{i}({\bm{X}},\lambda)=\sum_{j=1}^{r}\frac{\sigma_{j}^{2}u_{ij}^{2}}{\sigma_{j}^{2}+\lambda}$ (21)
Since $\bm{U}$} is an orthogonal matrix, for all $i\in[n]$,
$\sum_{j=1}^{r}u_{ij}^{2}\leq 1.$
Without loss of generality suppose $\sum_{j=1}^{r}u_{ij}^{2}=1$ since scaling up $u_{ij}$’s only increases the right-hand side of (21). By definition of $\zeta$, for all $i\in[n]$,
$\ell_{i}=\left\|{\bm{x}}_{i}\right\|_{2}^{2}=\sum_{j=1}^{r}\sigma_{j}^{2}u_{ij}^{2}\leq\zeta^{2}.$
Therefore by (21),
$h_{i}({\bm{X}},c\cdot\zeta^{2})\leq\sum_{j=1}^{r}\frac{\sigma_{j}^{2}u_{ij}^{2}}{\sigma_{j}^{2}+c\ell_{i}}$
Let $Y$ be a random variable that is equal to $\sigma_{j}^{2}$ with probability $u_{ij}^{2}$ and $\phi:{\mathbb{R}}_{\geq 0}\to{\mathbb{R}}_{\geq 0}$ be a function with $\phi(g)=\frac{g}{g+c\ell_{i}}$. Therefore
$\sum_{j=1}^{r}\frac{\sigma_{j}^{2}u_{ij}^{2}}{\sigma_{j}^{2}+c\ell_{i}}={\mathbb{E}}[\phi(Y)].$
Now by Jensen’s ienquality, since $\phi$ is a concave function, ${\mathbb{E}}[\phi(Y)]\leq\phi({\mathbb{E}}[Y])$. Therefore recalling that $\ell_{i}=\sum_{j=1}^{r}\sigma_{j}^{2}u_{ij}^{2}$, we have
$\sum_{j=1}^{r}\frac{\sigma_{j}^{2}u_{ij}^{2}}{\sigma_{j}^{2}+c\ell_{i}}\leq\phi({\mathbb{E}}[Y])=\frac{\ell_{i}}{\ell_{i}+c\ell_{i}}=\frac{1}{1+c}\,.$
Lemma A.2 ((Lemma 3.2 of miller1974unbalanced)).
Let
$\widehat{\boldsymbol{\beta}}=\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in\mathbb{R}^{k}}\left\|{\bm{y}}-{\bm{X}}{\bm{b}}\right\|_{2}^{2}~~\text{and}~~\widehat{\boldsymbol{\beta}}^{(-i)}=\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in\mathbb{R}^{k}}\left\|{\bm{y}}_{-i}-{\bm{X}}_{-i}{\bm{b}}\right\|_{2}^{2}.$
Then
$\widehat{\boldsymbol{\beta}}-\widehat{\boldsymbol{\beta}}^{(-i)}=\frac{({\bm{X}}^{\top}{\bm{X}})^{-1}{\bm{x}}_{i}(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}})}{1-h_{ii}},$
where $h_{ii}={\bm{x}}_{i}^{\top}({\bm{X}}^{\top}{\bm{X}})^{-1}{\bm{x}}_{i}$ is the leverage score for observation $i$.
Proof of Lemma˜B.1.
Let
[…]
One can observe that for any $\bm{b}$},
[…]
Therefore it immediately follows by Lemma˜A.2 that
${\bm{x}}_{i}^{\top}\boldsymbol{\beta}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}^{(-i)}=\frac{{\bm{x}}_{i}^{\top}({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{x}}_{i}(y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta})}{1-h_{i}({\bm{X}},\lambda)}=\frac{h_{i}({\bm{X}},\lambda)(y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta})}{1-h_{i}({\bm{X}},\lambda)}$
$\implies$ ${\bm{x}}_{i}^{\top}\boldsymbol{\beta}-y_{i}=(1-h_{i}({\bm{X}},\lambda))({\bm{x}}_{i}^{\top}\boldsymbol{\beta}^{(-i)}-y_{i})\,.$
Lemma A.3.
Let $\lambda^{(1)}\leq\lambda^{(2)}$ and for $j\in\{1,2\}$, ${\bm{v}}\in{\mathbb{R}}^{n}$, and
$\boldsymbol{\beta}_{\lambda^{(j)}}=\mathop{\mathrm{argmin}\,}\limits_{\boldsymbol{\beta}}\left\|{\bm{X}}\boldsymbol{\beta}-{\bm{v}}\right\|_{2}^{2}+\lambda^{(j)}\left\|\boldsymbol{\beta}\right\|_{2}^{2}.$
Then $\left\|{\bm{X}}\boldsymbol{\beta}_{\lambda^{(1)}}-{\bm{v}}\right\|_{2}^{2}\leq\left\|{\bm{X}}\boldsymbol{\beta}_{\lambda^{(2)}}-{\bm{v}}\right\|_{2}^{2}.$
Proof.
Let ${\bm{U}}{\boldsymbol{\Sigma}}{\bm{V}}^{\top}$ be the SVD of $\bm{X}$}. Then
${\bm{X}}({\bm{X}}^{\top}{\bm{X}}+\lambda^{(j)}{\bm{I}})^{-1}{\bm{X}}^{\top}={\bm{U}}{\boldsymbol{\Sigma}}({\boldsymbol{\Sigma}}^{\top}{\boldsymbol{\Sigma}}+\lambda^{(j)}{\bm{I}})^{-1}{\boldsymbol{\Sigma}}^{\top}{\bm{U}}^{\top}$
Therefore
$({\bm{I}}-{\bm{X}}({\bm{X}}^{\top}{\bm{X}}+\lambda^{(j)}{\bm{I}})^{-1}{\bm{X}}^{\top})^{2}={\bm{U}}({\bm{I}}-{\boldsymbol{\Sigma}}({\boldsymbol{\Sigma}}^{\top}{\boldsymbol{\Sigma}}+\lambda^{(j)}{\bm{I}})^{-1}{\boldsymbol{\Sigma}}^{\top})^{2}{\bm{U}}^{\top}$
Note ${\boldsymbol{\Sigma}}^{\top}{\boldsymbol{\Sigma}}+\lambda^{(1)}{\bm{I}}\preceq{\boldsymbol{\Sigma}}^{\top}{\boldsymbol{\Sigma}}+\lambda^{(2)}{\bm{I}}$. Therefore
${\bm{I}}-{\boldsymbol{\Sigma}}({\boldsymbol{\Sigma}}^{\top}{\boldsymbol{\Sigma}}+\lambda^{(1)}{\bm{I}})^{-1}{\boldsymbol{\Sigma}}^{\top}\preceq{\bm{I}}-{\boldsymbol{\Sigma}}({\boldsymbol{\Sigma}}^{\top}{\boldsymbol{\Sigma}}+\lambda^{(2)}{\bm{I}})^{-1}{\boldsymbol{\Sigma}}^{\top}.$
Now since both of these matrices are diagonal, we have
$({\bm{I}}-{\boldsymbol{\Sigma}}({\boldsymbol{\Sigma}}^{\top}{\boldsymbol{\Sigma}}+\lambda^{(1)}{\bm{I}})^{-1}{\boldsymbol{\Sigma}}^{\top})^{2}\preceq({\bm{I}}-{\boldsymbol{\Sigma}}({\boldsymbol{\Sigma}}^{\top}{\boldsymbol{\Sigma}}+\lambda^{(2)}{\bm{I}})^{-1}{\boldsymbol{\Sigma}}^{\top})^{2}.$
The result then follows by observing that
$\left\|{\bm{X}}\boldsymbol{\beta}^{(1)}-{\bm{y}}\right\|_{2}^{2}={\bm{y}}^{\top}({\bm{I}}-{\bm{X}}({\bm{X}}^{\top}{\bm{X}}+\lambda^{(j)}{\bm{I}})^{-1}{\bm{X}}^{\top})^{2}{\bm{y}}$
Appendix B Main Proofs
LOORA-HT depends on a set of $n$ leave-one-out regression coefficients $\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}$. Next lemma characterizes the difference between the full-sample regularized least-squares coefficient vector and its leave-one-out counterparts (see miller1974unbalanced for a simpler version for least-squares).
Lemma B.1.
Let $\lambda\geq 0$,
$\widehat{\boldsymbol{\beta}}_{\lambda}=\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in\mathbb{R}^{k}}\left\|{\bm{y}}-{\bm{X}}{\bm{b}}\right\|_{2}^{2}+\lambda\left\|{\bm{b}}\right\|_{2}^{2},~~\text{\ and\ }~~\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}=\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in\mathbb{R}^{k}}\left\|{\bm{y}}_{-i}-{\bm{X}}_{-i}{\bm{b}}\right\|_{2}^{2}+\lambda\left\|{\bm{b}}\right\|_{2}^{2}.$
Then,
$\widehat{\boldsymbol{\beta}}_{\lambda}-\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}=\frac{({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{x}}_{i}(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda})}{1-h_{\lambda ii}},$
and
$y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}=(1-h_{\lambda ii})(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}),$
where $h_{\lambda ii}={\bm{x}}_{i}^{\top}({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{x}}_{i}$. Moreover,
${\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}=\frac{{\bm{x}}_{i}^{\top}({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{X}}_{-i}^{\top}{\bm{y}}_{-i}}{1-h_{\lambda ii}}.$
Lemma˜B.1 basically gives us the intuition on the deviation that will happen by solving the left-one-out regression for the quantities such coefficients of the regression, value of errors, and estimations for each of the entries $i\in[n]$. The key point is that all the deviations depend on the value of the leverage score of the removed row, which further helps us characterize the robustness of our leave-one-out estimators to observation removal in inference time, which is a point of concern, especially in small sample settings where a single observation could have a large (close to $1$) leverage score that makes them affect quantities such as variance and confidence intervals to a large extent, which is not desirable. Hence, this is one of the first places that can motivate us to favor regularized regression adjustment methods, since as we will see in Lemma˜2, it will help us be able to uniformly decrease the leverage scores across all observations.
Lemma B.2.
Let
$g_{i}$ $=\frac{(r_{i}\mu_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})-{\bm{x}}_{i}^{\top}(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}_{-i}^{\top}({\bm{z}}{\bm{t}}/\boldsymbol{q})_{-i}}{q_{i}(1-\widetilde{h}_{\lambda ii})}\,.$
Under simple random assignment, the LOORA-HT estimator in Algorithm˜1 satisfies
$\widehat{\tau}_{\textup{LHT}}-\tau=\frac{1}{n}{\bm{z}}^{\top}\boldsymbol{g}.$
Proof.
First, notice that
$\widehat{\tau}_{\textup{LHT}}-\tau=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{z_{i}}{q_{i}}(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)})-(y^{(1)}_{i}-y^{(0)}_{i})\right)$.
Because $r_{i}\mu_{i}=(1-p_{i})y_{i}^{(1)}+p_{i}y_{i}^{(0)}$, it follows that
$\frac{z_{i}}{q_{i}}(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)})-(y^{(1)}_{i}-y^{(0)}_{i})=\frac{z_{i}}{q_{i}}(r_{i}\mu_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}).$
Algebraic manipulations yield $\widetilde{\bm{y}}-{\boldsymbol{\mu}}={\bm{z}}{\bm{t}}/\boldsymbol{q}$ and therefore,
$\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}=(\widetilde{{\bm{X}}}_{-i}^{\top}\widetilde{{\bm{X}}}_{-i}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}_{-i}^{\top}({\boldsymbol{\mu}}+{\bm{z}}{\bm{t}}/\boldsymbol{q})_{-i}\,.$
By Lemma˜B.1,
$\mu_{i}-\widetilde{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$ $=\mu_{i}-\widetilde{\bm{x}}_{i}^{\top}(\widetilde{{\bm{X}}}_{-i}^{\top}\widetilde{{\bm{X}}}_{-i}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}_{-i}^{\top}({\boldsymbol{\mu}}+{\bm{z}}{\bm{t}}/\boldsymbol{q})_{-i}$
$=\mu_{i}-\widetilde{\bm{x}}_{i}^{\top}(\widetilde{{\bm{X}}}_{-i}^{\top}\widetilde{{\bm{X}}}_{-i}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}_{-i}^{\top}{\boldsymbol{\mu}}_{-i}-\widetilde{\bm{x}}_{i}^{\top}(\widetilde{{\bm{X}}}_{-i}^{\top}\widetilde{{\bm{X}}}_{-i}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}_{-i}^{\top}({\bm{z}}{\bm{t}}/\boldsymbol{q})_{-i}$
$=\frac{\mu_{i}-\widetilde{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}-\widetilde{\bm{x}}_{i}^{\top}(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}_{-i}^{\top}({\bm{z}}{\bm{t}}/\boldsymbol{q})_{-i}}{1-\widetilde{h}_{\lambda ii}},$
where $\boldsymbol{\beta}_{\lambda}$ is defined as in Equation˜6. Therefore,
$\frac{z_{i}}{q_{i}}(r_{i}\mu_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)})$ $=\frac{z_{i}}{q_{i}}r_{i}(\mu_{i}-\widetilde{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)})$
$=z_{i}\frac{(r_{i}\mu_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})-{\bm{x}}_{i}^{\top}(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}_{-i}^{\top}({\bm{z}}{\bm{t}}/\boldsymbol{q})_{-i}}{q_{i}(1-\widetilde{h}_{\lambda ii})}.$
B.1 Omitted Proofs of Section˜3
Proof of Theorem˜1.
Because $z_{i}$ is independent of $\widehat{\boldsymbol{\beta}}^{(-i)}$,
${\mathbb{E}}\left[\frac{z_{i}}{q_{i}}(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}^{(-i)})\right]$ $={\mathbb{E}}[y^{(1)}_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}^{(-i)}|z_{i}=1]-{\mathbb{E}}[y^{(0)}_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}^{(-i)}|z_{i}=-1]$
$=y^{(1)}_{i}-y^{(0)}_{i}\,$.
As a result, $\widehat{\tau}_{\textup{LHT}}$ is unbiased. By Lemma˜B.2
${\mathbb{E}}[(\widehat{\tau}_{\textup{LHT}}-\tau)^{2}]=\frac{1}{n^{2}}{\mathbb{E}}[{\bm{z}}^{\top}\boldsymbol{g}\boldsymbol{g}^{\top}{\bm{z}}].$
By Lemma˜B.2, we have $\widehat{\tau}-\tau=\frac{1}{n}{\bm{z}}^{\top}\boldsymbol{g}$. Therefore
${\mathbb{E}}[(\widehat{\tau}-\tau)^{2}]=\frac{1}{n^{2}}{\mathbb{E}}[{\bm{z}}^{\top}\boldsymbol{g}\boldsymbol{g}^{\top}{\bm{z}}].$
To calculate ${\mathbb{E}}[{\bm{z}}^{\top}\boldsymbol{g}\boldsymbol{g}^{\top}{\bm{z}}]$, first notice that $z_{i}^{2}=1$ and ${\mathbb{E}}[1/q_{i}^{2}]=1/r_{i}^{2}$. Therefore,
${\mathbb{E}}\left[\left(\frac{z_{i}(r_{i}\mu_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})}{q_{i}(1-\widetilde{h}_{\lambda ii})}\right)^{2}\right]$ $=\frac{(r_{i}\mu_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})^{2}}{r_{i}^{2}(1-\widetilde{h}_{\lambda ii})^{2}}=\frac{(\mu_{i}-\widetilde{{\bm{x}}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})^{2}}{(1-\widetilde{h}_{\lambda ii})^{2}}.$
Because $z_{i}/q_{i}$ and $z_{j}/q_{j}$ are independent for $i\neq j$, and ${\mathbb{E}}[z_{i}/q_{i}]=0$, then
${\mathbb{E}}\Bigg[\frac{z_{i}(r_{i}\mu_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})}{q_{i}(1-\widetilde{h}_{\lambda ii})}$ $\frac{z_{i}{\bm{x}}_{i}^{\top}(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}_{-i}^{\top}({\bm{z}}{\bm{t}}/\boldsymbol{q})_{-i}}{q_{i}(1-\widetilde{h}_{\lambda ii})}\Bigg]$
$={\mathbb{E}}\Bigg[\frac{(r_{i}\mu_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})}{q_{i}^{2}}\Bigg]\frac{{\bm{x}}_{i}^{\top}(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}_{-i}^{\top}}{(1-\widetilde{h}_{\lambda ii})^{2}}\,{\mathbb{E}}[({\bm{z}}{\bm{t}}/\boldsymbol{q})_{-i}]=0.$
Moreover for $i\neq j$,
${\mathbb{E}}\Bigg[\frac{z_{i}(r_{i}\mu_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})}{q_{i}(1-\widetilde{h}_{\lambda ii})}$ $\frac{z_{j}{\bm{x}}_{i}^{\top}(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}_{-j}^{\top}({\bm{z}}{\bm{t}}/\boldsymbol{q})_{-j}}{q_{j}(1-\widetilde{h}_{\lambda jj})}\Bigg]$
$={\mathbb{E}}\Bigg[\frac{z_{j}}{q_{j}}\Bigg]{\mathbb{E}}\Bigg[\frac{z_{i}(r_{i}\mu_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})}{q_{i}(1-\widetilde{h}_{\lambda ii})}\frac{{\bm{x}}_{i}^{\top}(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}_{-j}^{\top}({\bm{z}}{\bm{t}}/\boldsymbol{q})_{-j}}{(1-\widetilde{h}_{\lambda jj})}\Bigg]=0.$
Using ${\mathbb{E}}[1/q_{i}^{2}]=1/r_{i}^{2}$ and independent between $z_{i}/q_{i}$ and $z_{j}/q_{j}$ for $i\neq j$, we obtain
${\mathbb{E}}\left[\left(\frac{z_{i}{\bm{x}}_{i}^{\top}(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}_{-i}^{\top}({\bm{z}}{\bm{t}}/\boldsymbol{q})_{-i}}{q_{i}(1-\widetilde{h}_{\lambda ii})}\right)^{2}\right]$ $=\frac{1}{(1-\widetilde{h}_{\lambda ii})^{2}}{\mathbb{E}}\Bigg[\Bigg(\sum_{j\neq i}\widetilde{h}_{\lambda ij}t_{j}z_{j}/q_{j}\Bigg)^{2}\Bigg]$
$=\sum_{j\neq i}\frac{\widetilde{h}_{\lambda ij}^{2}t_{j}^{2}}{r_{j}^{2}(1-\widetilde{h}_{\lambda ii})^{2}}.$
Because $z_{i}/q_{i}$ and $z_{j}/q_{j}$ are independent for $i\neq j$ and ${\mathbb{E}}[z_{i}/q_{i}]=0$,
${\mathbb{E}}\left[\frac{z_{i}(r_{i}\mu_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta})}{q_{i}(1-\widetilde{h}_{\lambda ii})}\frac{z_{j}(r_{j}\mu_{j}-{\bm{x}}_{j}^{\top}\boldsymbol{\beta})}{q_{j}(1-\widetilde{h}_{\lambda jj})}\right]=0.$
Finally,
${\mathbb{E}}$ $\left[\frac{z_{i}{\bm{x}}_{i}^{\top}(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}_{-i}^{\top}({\bm{z}}{\bm{t}}/\boldsymbol{q})_{-i}}{q_{i}(1-\widetilde{h}_{\lambda ii})}\frac{z_{j}{\bm{x}}_{j}^{\top}(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}_{-j}^{\top}({\bm{z}}{\bm{t}}/\boldsymbol{q})_{-j}}{q_{j}(1-\widetilde{h}_{\lambda jj})}\right]$
$={\mathbb{E}}\left[\frac{z_{i}r_{i}\widetilde{\bm{x}}_{i}^{\top}(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}_{-i}^{\top}({\bm{z}}{\bm{t}}/\boldsymbol{q})_{-i}}{q_{i}(1-\widetilde{h}_{\lambda ii})}\frac{z_{j}r_{j}\widetilde{\bm{x}}_{j}^{\top}(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}_{-j}^{\top}({\bm{z}}{\bm{t}}/\boldsymbol{q})_{-j}}{q_{j}(1-\widetilde{h}_{\lambda jj})}\right]$
$={\mathbb{E}}\left[\frac{r_{i}r_{j}\widetilde{h}_{\lambda ij}^{2}t_{i}t_{j}}{q_{i}^{2}q_{j}^{2}(1-\widetilde{h}_{\lambda ii})(1-\widetilde{h}_{\lambda jj})}\right]$
$=\frac{\widetilde{h}_{\lambda ij}^{2}t_{i}t_{j}}{r_{i}r_{j}(1-\widetilde{h}_{\lambda ii})(1-\widetilde{h}_{\lambda jj})}.$
Combining all the above, we have
$\frac{1}{n^{2}}$ $\left(\sum_{i=1}^{n}\frac{(\mu_{i}-\widetilde{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})^{2}}{(1-\widetilde{h}_{\lambda ii})^{2}}+\sum_{i=1}^{n}\sum_{j\neq i}\frac{(\widetilde{h}_{\lambda ij}t_{j})^{2}}{r_{j}^{2}(1-\widetilde{h}_{\lambda ii})^{2}}+\frac{\widetilde{h}_{\lambda ij}^{2}~t_{i}t_{j}}{r_{i}r_{j}(1-\widetilde{h}_{\lambda ii})(1-\widetilde{h}_{\lambda jj})}\right)$
$=\frac{1}{n^{2}}\left(\sum_{i=1}^{n}\frac{(\mu_{i}-\widetilde{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})^{2}}{(1-\widetilde{h}_{\lambda ii})^{2}}+\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}\widetilde{h}_{\lambda ij}^{2}\left(\frac{t_{j}}{r_{j}(1-\widetilde{h}_{\lambda ii})}+\frac{t_{i}}{r_{i}(1-\widetilde{h}_{\lambda jj})}\right)^{2}\right).$
AA about here.
Proof of Lemma˜3.
First, observe that
$\boldsymbol{\beta}^{*}_{\lambda}=\lim_{n\to\infty}(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}^{\top}{\boldsymbol{\mu}}.$
Under Assumption˜2, $(n^{-1}{\bm{X}}^{\top}{\bm{X}})\to\Sigma_{X}$, where $\Sigma_{X}$ is bounded and positive definite. Moreover, by Assumption˜3, since $m=\inf_{i\in\mathbb{N}}\min\{p_{i},1-p_{i}\}>0$, we have
$n^{-1}\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}=n^{-1}{\bm{X}}^{\top}{\bm{R}}^{-2}{\bm{X}}\to\Sigma_{\widetilde{X}}$,
where $\Sigma_{\widetilde{X}}$ is bounded and invertible. Hence,
$\left\|(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\right\|_{2}=O(n^{-1}).$ (22)
Furthermore,
$\widetilde{{\bm{X}}}^{\top}{\boldsymbol{\mu}}=\sum_{i=1}^{n}{\bm{x}}_{i}\left(\frac{y_{i}^{(1)}}{p_{i}}+\frac{y_{i}^{(0)}}{1-p_{i}}\right),$
and by Assumptions˜2 and 3, $\left\|\widetilde{{\bm{X}}}^{\top}{\boldsymbol{\mu}}\right\|_{2}=O(n)$. It follows that the limit $(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}^{\top}{\boldsymbol{\mu}}$ exists and that $\boldsymbol{\beta}^{*}_{\lambda}$ is finite.
Step 1: Consistency of the full-sample estimator. Define
$\widehat{\boldsymbol{\beta}}_{\lambda}=(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{y}}}$.
By construction, $\widetilde{{\bm{y}}}={\boldsymbol{\mu}}+{\bm{z}}{\bm{t}}/\boldsymbol{q}$, where ${\mathbb{E}}[z_{i}t_{i}/q_{i}]=0$ for all $i$. Because ${\bm{x}}_{i}$, $t_{i}$, and $1/p_{i},1/(1-p_{i})$ are uniformly bounded, the weak law of large numbers implies
$\frac{1}{n}\widetilde{{\bm{X}}}^{\top}({\bm{z}}{\bm{t}}/\boldsymbol{q})\xrightarrow{p}\boldsymbol{0}.$
By (22) and Slutsky’s theorem,
$(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{y}}}\xrightarrow{p}(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}^{\top}{\boldsymbol{\mu}},$
and hence $\widehat{\boldsymbol{\beta}}_{\lambda}\xrightarrow{p}\boldsymbol{\beta}^{*}_{\lambda}$.
Step 2: Asymptotic equivalence of leave-one-out and full-sample estimators. By Lemma˜B.1,
$\widehat{\boldsymbol{\beta}}_{\lambda}-\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}=\frac{(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{x}}}_{i}(\widetilde{y}_{i}-\widetilde{{\bm{x}}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda})}{1-\widetilde{h}_{\lambda ii}},$ (23)
where $\widetilde{h}_{\lambda ii}=\widetilde{{\bm{x}}}_{i}^{\top}(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{x}}}_{i}$. From Assumptions˜1 and 3 and (22),
$\widetilde{h}_{\lambda ii}\leq\|\widetilde{{\bm{x}}}_{i}\|_{2}^{2}\,\left\|(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\right\|_{2}=O(n^{-1}),$
so $\widetilde{h}_{\lambda ii}\to 0$ uniformly in $i$, and the denominator in (23) converges to $1$.
Next, by Assumption˜1 and the inequalities $\|{\bm{y}}^{(1)}\|_{\infty}\leq(nL)^{1/4}$ and $\|{\bm{y}}^{(0)}\|_{\infty}\leq(nL)^{1/4}$, the responses are uniformly bounded in $O(n^{1/4})$. Because $\|(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\|_{2}=O(n^{-1})$ and $\|\widetilde{{\bm{x}}}_{i}\|_{2}=O(1)$, it follows from (23) that
$\left\|\widehat{\boldsymbol{\beta}}_{\lambda}-\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}\right\|_{2}=O(n^{-3/4}).$
Hence $\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}\xrightarrow{p}\widehat{\boldsymbol{\beta}}_{\lambda}$ for each $i$. Combining Steps 1 and 2,
$\left\|\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}-\boldsymbol{\beta}^{*}_{\lambda}\right\|_{2}\leq\left\|\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}-\widehat{\boldsymbol{\beta}}_{\lambda}\right\|_{2}+\left\|\widehat{\boldsymbol{\beta}}_{\lambda}-\boldsymbol{\beta}^{*}_{\lambda}\right\|_{2}\xrightarrow{p}0,$
which completes the proof. ∎
Proof of Theorem˜4.
By Lemma˜3, for all $i\in\mathbb{N}$, $\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}\xrightarrow{p}\boldsymbol{\beta}^{*}_{\lambda}$. Hence, by Slutsky’s theorem,
${\widehat{\tau}}_{\textup{LHT}}=\frac{1}{n}\sum_{i=1}^{n}\frac{d_{i}\big(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}\big)}{p_{i}}-\frac{1}{n}\sum_{i=1}^{n}\frac{(1-d_{i})\big(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}\big)}{1-p_{i}}\;\xrightarrow{d}\;{\widehat{\tau}}_{\textup{LHT}}^{*},$
where
${\widehat{\tau}}_{\textup{LHT}}^{*}:=\frac{1}{n}\sum_{i=1}^{n}\frac{d_{i}\big(y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}^{*}_{\lambda}\big)}{p_{i}}-\frac{1}{n}\sum_{i=1}^{n}\frac{(1-d_{i})\big(y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}^{*}_{\lambda}\big)}{1-p_{i}}.$
Using independence of $d_{i}$’s under simple random assignment together with ${\mathbb{E}}[d_{i}-p_{i}]=0$, we obtain
${\rm Var}\big({\widehat{\tau}}_{\textup{LHT}}^{*}\big)=\frac{1}{n^{2}}\,\big\|\widetilde{{\bm{X}}}\boldsymbol{\beta}_{\lambda}^{*}-{\boldsymbol{\mu}}\big\|_{2}^{2}.$
By Assumptions˜1 and 3, the summands of ${\widehat{\tau}}_{\textup{LHT}}^{*}$ are independent, mean-zero after centering at $\tau$, and uniformly bounded in absolute value. Hence the Lindeberg condition holds, and by the Lindeberg–Feller CLT (e.g., Theorems 5.2 and 5.10 of white),
$\sqrt{n}\,\big({\widehat{\tau}}_{\textup{LHT}}^{*}-\tau\big)\;\xrightarrow{d}\;\mathcal{N}\!\left(0,\;\lim_{n\to\infty}\frac{1}{n}\,\big\|\widetilde{{\bm{X}}}\boldsymbol{\beta}_{\lambda}^{*}-{\boldsymbol{\mu}}\big\|_{2}^{2}\right).$
It remains to show
$\sqrt{n}\,\big({\widehat{\tau}}_{\textup{LHT}}-{\widehat{\tau}}_{\textup{LHT}}^{*}\big)\;\xrightarrow{p}\;0,$
so that the same limit law transfers to ${\widehat{\tau}}_{\textup{LHT}}$. Define
$a_{i}:=\frac{d_{i}}{p_{i}}-\frac{1-d_{i}}{1-p_{i}}=\frac{d_{i}-p_{i}}{p_{i}(1-p_{i})},\qquad\Delta_{i}:=\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}-\boldsymbol{\beta}^{*}_{\lambda}.$
Then
${\widehat{\tau}}_{\textup{LHT}}-{\widehat{\tau}}_{\textup{LHT}}^{*}=-\,\frac{1}{n}\sum_{i=1}^{n}a_{i}\,{\bm{x}}_{i}^{\top}\Delta_{i}.$
Split $\Delta_{i}=(\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}-\widehat{\boldsymbol{\beta}}_{\lambda})+(\widehat{\boldsymbol{\beta}}_{\lambda}-\boldsymbol{\beta}^{*}_{\lambda})$. For the first part, by the leave-one-out identity (cf. (23)) and the bounds already established in the previous lemma,
$\max_{i\leq n}\big\|\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}-\widehat{\boldsymbol{\beta}}_{\lambda}\big\|_{2}=O_{p}\!\big(n^{-3/4}\big).$
Since $\|a_{i}\|_{\infty}\leq m^{-1}(1-m)^{-1}$ and $\|{\bm{x}}_{i}\|_{2}=O(1)$ by Assumptions˜1 and 3, we get
$\sqrt{n}\left|\frac{1}{n}\sum_{i=1}^{n}a_{i}\,{\bm{x}}_{i}^{\top}\big(\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}-\widehat{\boldsymbol{\beta}}_{\lambda}\big)\right|\leq\sqrt{n}\,\frac{1}{n}\sum_{i=1}^{n}O(1)\cdot O_{p}\!\big(n^{-3/4}\big)=O_{p}\!\big(n^{-1/4}\big)\;\to\;0.$
For the second part,
$\sqrt{n}\left|\frac{1}{n}\sum_{i=1}^{n}a_{i}\,{\bm{x}}_{i}^{\top}\big(\widehat{\boldsymbol{\beta}}_{\lambda}-\boldsymbol{\beta}^{*}_{\lambda}\big)\right|=\left\|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}a_{i}\,{\bm{x}}_{i}\right\|_{2}\cdot\big\|\widehat{\boldsymbol{\beta}}_{\lambda}-\boldsymbol{\beta}^{*}_{\lambda}\big\|_{2}.$
By a standard CLT for independent summands with bounded second moments,
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}a_{i}{\bm{x}}_{i}=O_{p}(1).$
Moreover, under Assumptions˜1 and 2, the ridge estimator with fixed $\lambda\geq 0$ satisfies $\|\widehat{\boldsymbol{\beta}}_{\lambda}-\boldsymbol{\beta}^{*}_{\lambda}\|_{2}=O_{p}(n^{-1/2})$. Hence the product is $O_{p}(n^{-1/2})\to 0$. Combining the two parts yields $\sqrt{n}\,({\widehat{\tau}}_{\textup{LHT}}-{\widehat{\tau}}_{\textup{LHT}}^{*})\xrightarrow{p}0$. Therefore, by the asymptotic equivalence lemma, e.g., (rao73, pg. 122), the limit distribution for $\sqrt{n}({\widehat{\tau}}_{\textup{LHT}}-\tau)$ coincides with that of $\sqrt{n}({\widehat{\tau}}_{\textup{LHT}}^{*}-\tau)$, namely a mean-zero Gaussian with variance $\lim_{n\to\infty}\frac{1}{n}\|\widetilde{{\bm{X}}}\boldsymbol{\beta}_{\lambda}^{*}-{\boldsymbol{\mu}}\|_{2}^{2}$. ∎
B.2 Omitted Proofs of Section˜4
Proof of Theorem˜5.
We first show that the estimator is unbiased. Let
${\boldsymbol{\Lambda}}_{\lambda}^{(-i)}=({\bm{X}}_{-i}^{\top}{\bm{X}}_{-i}+\lambda{\bm{I}})^{-1}.$
By law of total expectation, we have
${\mathbb{E}}\left[v_{i}z_{i}(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)})\right]$
$=\frac{n_{T}}{n}{\mathbb{E}}\left[\frac{1}{n_{T}}(y^{(1)}_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)})|z_{i}=+1\right]-\frac{n_{C}}{n}{\mathbb{E}}\left[\frac{1}{n_{C}}(y^{(0)}_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)})|z_{i}=-1\right]$
$=\frac{1}{n}(y^{(1)}_{i}-y^{(0)}_{i})-\frac{n_{C}n_{T}(n-1)}{n^{2}}{\mathbb{E}}\left[{\bm{x}}_{i}^{\top}{\boldsymbol{\Lambda}}_{\lambda}^{(-i)}{\bm{X}}_{-i}^{\top}({\bm{f}}^{(T)}{\bm{y}})_{-i}|z_{i}=+1\right]$
$+\frac{n_{C}n_{T}(n-1)}{n^{2}}{\mathbb{E}}\left[{\bm{x}}_{i}^{\top}{\boldsymbol{\Lambda}}_{\lambda}^{(-i)}{\bm{X}}_{-i}^{\top}({\bm{f}}^{(C)}{\bm{y}})_{-i}|z_{i}=-1\right]\,.$
Now by linearity of expectation,
${\mathbb{E}}\left[{\bm{x}}_{i}^{\top}{\boldsymbol{\Lambda}}_{\lambda}^{(-i)}{\bm{X}}_{-i}^{\top}({\bm{f}}^{(T)}{\bm{y}})_{-i}|z_{i}=+1\right]=({\bm{x}}_{i}^{\top}{\boldsymbol{\Lambda}}_{\lambda}^{(-i)}{\bm{X}}_{-i}^{\top}){\mathbb{E}}\left[({\bm{f}}^{(T)}{\bm{y}})_{-i}|z_{i}=+1\right]\,.$
Moreover
${\mathbb{E}}\left[({\bm{f}}^{(T)}{\bm{y}})_{-i}|z_{i}=+1\right]$ $=\frac{n_{T}-1}{n-1}\cdot\frac{{\bm{y}}^{(1)}_{-i}}{n_{T}(n_{T}-1)}+\frac{n_{C}}{n-1}\cdot\frac{{\bm{y}}^{(0)}_{-i}}{n_{C}^{2}}=\frac{({\bm{y}}^{(1)}/n_{T}+{\bm{y}}^{(0)}/n_{C})_{-i}}{n-1}\,.$
Similarly
${\mathbb{E}}\left[({\bm{f}}^{(C)}{\bm{y}})_{-i}|z_{i}=-1\right]=\frac{({\bm{y}}^{(1)}/n_{T}+{\bm{y}}^{(0)}/n_{C})_{-i}}{n-1}\,.$
Therefore
${\mathbb{E}}\left[v_{i}z_{i}(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)})\right]=\frac{1}{n}(y^{(1)}_{i}-y^{(0)}_{i})$,
and LOORA-DM is an unbiased estimator. We define $\widetilde{{\bm{t}}}^{(-i)}$ and ${\bm{v}}^{(-i)}$ as the following.
$\widetilde{{\bm{t}}}^{(-i)}=\begin{cases}\widetilde{{\bm{t}}}^{(1)}&\text{if }z_{i}=1,\\ \widetilde{{\bm{t}}}^{(0)}&\text{if }z_{i}=-1.\end{cases}$ (24)
[…]
Algebraic manipulations yield
$\widetilde{{\bm{y}}}^{(-i)}=\frac{\widetilde{{\boldsymbol{\mu}}}}{n}+\frac{{\bm{v}}^{(-i)}{\bm{z}}\widetilde{{\bm{t}}}^{(-i)}}{n}.$ (26)
Moreover, irresepective of the assignment of unit $i$,
$v_{i}z_{i}(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)})-\frac{1}{n}(y^{(1)}_{i}-y^{(0)}_{i})=v_{i}z_{i}(\widetilde{{\boldsymbol{\mu}}}_{i}/n-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)})\,.$
Therefore
$\mathrm{Var}(\widehat{\tau})\;=\;{\mathbb{E}}\bigl[(\widehat{\tau}-\tau)^{2}\bigr]\;=\;{\mathbb{E}}\Biggl[\Bigl(\sum_{i=1}^{n}v_{i}\,z_{i}\,\Bigl(\tfrac{\widetilde{\boldsymbol{\mu}}_{i}}{n}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}^{(-i)}\Bigr)\Bigr)^{2}\Biggr].$ (27)
By (26),
$\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}\;=\;{\boldsymbol{\Lambda}}_{\lambda}^{(-i)}\,{\bm{X}}_{-i}^{\top}\,\left(\tfrac{\widetilde{\boldsymbol{\mu}}}{n}+\frac{{\bm{v}}^{(-i)}{\bm{z}}\widetilde{\bm{t}}^{(-i)}}{n}\right)_{-i}.$
We denote $\boldsymbol{\Lambda}_{\lambda}=({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}$. Substituting these into (27) and a short calculation yields
[…]
$={\mathbb{E}}\Biggl[\Biggl(\sum_{i=1}^{n}v_{i}\,z_{i}\,\Bigl\{\frac{\widetilde{\boldsymbol{\mu}}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{n\,(1-h_{\lambda ii})}-\frac{{\bm{x}}_{i}^{\top}\,\boldsymbol{\Lambda}_{\lambda}\,{\bm{X}}_{-i}^{\top}}{(1-h_{\lambda ii})}\,\left(\frac{{\bm{v}}^{(-i)}{\bm{z}}\widetilde{\bm{t}}^{(-i)}}{n}\right)_{-i}\Bigr\}\Biggr)^{2}\Biggr],$ (29)
where the second equality follows from Lemmas˜A.2 and B.1. For a more compact notation, we denote $\overline{h}_{i}:=(1-h_{\lambda ii})$ for the rest of this proof. We decompose (29) into three components and compute the corresponding expectations separately.
  1. 1.
    $T_{1}={\mathbb{E}}\left[\sum_{i,j\in[n]}\frac{v_{i}\,z_{i}\,v_{j}\,z_{j}\,}{n^{2}}\frac{\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{\overline{h}_{i}}\,\frac{\widetilde{\mu}_{j}-{\bm{x}}_{j}^{\top}\boldsymbol{\beta}_{\lambda}}{\overline{h}_{j}}\right]$.
  2. 2.
    $T_{2}=-2\,{\mathbb{E}}\left[\sum_{i,j\in[n]}\frac{v_{i}\,z_{i}\,v_{j}\,z_{j}\,}{n^{2}}\frac{\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{\overline{h}_{i}}\;\left({\bm{x}}_{j}^{\top}\,\boldsymbol{\Lambda}_{\lambda}\,{\bm{X}}_{-j}^{\top}\,({\bm{v}}^{(-j)}{\bm{z}}\widetilde{\bm{t}}^{(-j)})_{-j}\right)\right]$.
  3. 3.
    $T_{3}={\mathbb{E}}\left[\sum\limits_{i,j\in[n]}\!\!\frac{v_{i}\,z_{i}\,v_{j}\,z_{j}\,}{n^{2}\overline{h}_{i}\overline{h}_{j}}\!\!\left({\bm{x}}_{i}^{\top}\,\boldsymbol{\Lambda}_{\lambda}\,{\bm{X}}_{-i}^{\top}\,({\bm{v}}^{(-i)}{\bm{z}}\widetilde{\bm{t}}^{(-i)})_{-i}\right)\!\!\left({\bm{x}}_{j}^{\top}\,\boldsymbol{\Lambda}_{\lambda}\,{\bm{X}}_{-j}^{\top}\,({\bm{v}}^{(-j)}{\bm{z}}\widetilde{\bm{t}}^{(-j)})_{-j}\right)\right]$.
One can easily observe that
${\rm Var}(\widehat{\tau})\;=\;T_{1}+T_{2}+T_{3}.$
We start by calculating $T_{1}$. Note that
$T_{1}={\mathbb{E}}\Biggl[\Biggl(\sum_{i=1}^{n}\frac{v_{i}\,z_{i}\,}{n}\frac{\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{\overline{h}_{i}}\Biggr)^{2}\Biggr].$
We denote $R_{i}:=\frac{\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{\overline{h}_{i}}$. Since $\widetilde{{\boldsymbol{\mu}}}$ and $\boldsymbol{\beta}_{\lambda}$ are fixed vectors and do not depend on the treatment assignment vector $\bm{z}$}, by linearity of expectation,
$T_{1}=\frac{1}{n^{2}}{\mathbb{E}}\Biggl[\Biggl(\sum_{i=1}^{n}v_{i}z_{i}R_{i}\Biggr)^{2}\Biggr]=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}R_{i}R_{j}{\mathbb{E}}[v_{i}z_{i}v_{j}z_{j}].$
For $i=j$, ${\mathbb{E}}[(v_{i}z_{i})^{2}]={\mathbb{E}}[v_{i}^{2}]$. Therefore by defnition,
${\mathbb{E}}[v_{i}z_{i}v_{j}z_{j}]={\mathbb{E}}[v_{i}^{2}]=\frac{n_{T}}{n}\left(\frac{1}{n_{T}}\right)^{2}+\frac{n_{C}}{n}\left(\frac{1}{n_{C}}\right)^{2}=\frac{1}{nn_{T}}+\frac{1}{nn_{C}}=\frac{n_{C}+n_{T}}{nn_{T}n_{C}}=\frac{1}{n_{T}n_{C}}.$
For $i\neq j$, we can write ${\mathbb{E}}[v_{i}z_{i}v_{j}z_{j}]$ by considering the joint assignment of units $i$ and $j$.
${\mathbb{E}}[v_{i}z_{i}v_{j}z_{j}]$ $=\mathbb{P}(d_{i}=1,d_{j}=1)\left(\frac{1}{n_{T}}\cdot\frac{1}{n_{T}}\right)+\mathbb{P}(d_{i}=1,d_{j}=0)\left(\frac{1}{n_{T}}\cdot\frac{-1}{n_{C}}\right)$
$+\mathbb{P}(d_{i}=0,d_{j}=1)\left(\frac{-1}{n_{C}}\cdot\frac{1}{n_{T}}\right)+\mathbb{P}(d_{i}=0,d_{j}=0)\left(\frac{-1}{n_{C}}\cdot\frac{-1}{n_{C}}\right)$
$=\frac{n_{T}(n_{T}-1)}{n(n-1)}\frac{1}{n_{T}^{2}}-\frac{n_{T}n_{C}}{n(n-1)}\frac{1}{n_{T}n_{C}}-\frac{n_{C}n_{T}}{n(n-1)}\frac{1}{n_{C}n_{T}}+\frac{n_{C}(n_{C}-1)}{n(n-1)}\frac{1}{n_{C}^{2}}$
$=-\frac{1}{(n-1)n_{T}n_{C}}.$
Substituting these into the expression for $T_{1}$, we have
$T_{1}$ $=\frac{1}{n^{2}}\left(\sum_{i=1}^{n}R_{i}^{2}{\mathbb{E}}[v_{i}^{2}]+\sum_{i\neq j}R_{i}R_{j}{\mathbb{E}}[v_{i}z_{i}v_{j}z_{j}]\right)$
$=\frac{1}{n^{2}n_{T}n_{C}}\left(\sum_{i=1}^{n}R_{i}^{2}-\frac{1}{n-1}\sum_{i,j\in[n]:i\neq j}R_{i}R_{j}\right).$
Using the identity $\sum_{i,j\in[n]:i\neq j}R_{i}R_{j}=\left(\sum_{i\in[n]}R_{i}\right)^{2}-\sum_{i\in[n]}R_{i}^{2}$, simple calculations yield
$T_{1}$ $=\frac{1}{n(n-1)n_{T}n_{C}}\sum_{i=1}^{n}\left(R_{i}-\frac{1}{n}\sum_{j=1}^{n}R_{j}\right)^{2}$
$=\frac{1}{n(n-1)n_{T}n_{C}}\sum_{i=1}^{n}\left(\frac{\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{(1-h_{\lambda ii})}-\frac{1}{n}\sum_{j=1}^{n}\frac{\widetilde{\mu}_{j}-{\bm{x}}_{j}^{\top}\boldsymbol{\beta}_{\lambda}}{(1-h_{\lambda jj})}\right)^{2}.$
Recall that $\boldsymbol{H}_{\lambda}={\bm{X}}^{\top}\boldsymbol{\Lambda}_{\lambda}{\bm{X}}$. We have
$T_{2}$ $=-2\,{\mathbb{E}}\left[\Biggl(\sum_{i=1}^{n}\frac{v_{i}z_{i}R_{i}}{n}\Biggr)\Biggl(\sum_{j=1}^{n}\frac{v_{j}z_{j}}{n\overline{h}_{j}}{\bm{x}}_{j}^{\top}{\boldsymbol{\Lambda}}_{\lambda}{\bm{X}}_{-j}^{\top}({\bm{v}}^{(-j)}{\bm{z}}\widetilde{{\bm{t}}}^{(-j)})_{-j}\Biggr)\right]$
$=\frac{-2}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}\frac{R_{i}}{\overline{h}_{j}}\sum_{k\in[n]:k\neq j}h_{\lambda jk}\,{\mathbb{E}}\left[v_{i}z_{i}v_{j}z_{j}v_{k}^{(-j)}z_{k}\widetilde{t}_{k}^{(-j)}\right]$
We denote ${\mathbb{E}}\left[v_{i}z_{i}v_{j}z_{j}v_{k}^{(-j)}z_{k}\widetilde{t}_{k}^{(-j)}\right]$. One can easily check that $E_{iik}=0$ for $i\neq k$. Therefore we consider two cases: 1) $i=k$, and 2) $i\neq k$, $i\neq j$—note that the above expression guarantees $j\neq k$. We have
$E_{iji}={\mathbb{E}}\left[v_{i}v^{(-j)}_{i}v_{j}z_{j}\widetilde{t}_{i}^{(-j)}\right]=\frac{\widetilde{\mu}_{i}}{(n-1)n_{T}n_{C}}.$ (30)
For $i\neq k$, $i\neq j$, considering the joint assignments of $i,j,k$ yields,
$E_{ijk}={\mathbb{E}}\left[v_{i}z_{i}v_{j}z_{j}v_{k}^{(-j)}z_{k}\widetilde{t}_{k}^{(-j)}\right]=\frac{-\widetilde{\mu}_{k}}{(n-1)(n-2)n_{T}n_{C}}.$
We first consider the terms in $T_{2}$ with $i=k$. Note that in this case, since $j\neq k$, also $i\neq j$. Therefore by (30), and some calculations, the sum of those terms are equal to
$T_{2,A}:=\frac{-2}{n^{2}(n-1)n_{T}n_{C}}\sum_{i,j\in[n]:i\neq j}\frac{R_{i}h_{\lambda ij}\widetilde{\mu}_{i}}{\overline{h}_{j}}.$
Similar calculations reveal that the terms corresponding to the case $i\neq k$, $i\neq j$ add up to
$T_{2,B}=\frac{2}{n^{2}(n-1)(n-2)n_{T}n_{C}}\sum_{i,j\in[n]:i\neq j}\sum_{k\in[n]:k\neq i,j}\frac{R_{i}h_{\lambda jk}\widetilde{\mu}_{k}}{\overline{h}_{j}}.$
Therefore
$T_{2}=T_{2,A}+T_{2,B}=\frac{-2}{n^{2}(n-1)n_{T}n_{C}}\left(\sum_{i,j\in[n]:i\neq j}\frac{R_{i}h_{\lambda ij}\widetilde{\mu}_{i}}{\overline{h}_{j}}-\frac{1}{n-2}\sum_{k\in[n]:k\neq i,j}\frac{R_{i}h_{\lambda jk}\widetilde{\mu}_{k}}{\overline{h}_{j}}\right).$
Finally for $T_{3}$, following the same proof strategy as $T_{1},T_{2}$, and calculating the expectation over the joint assignment of four distinct units $(i,j,k,l)$ for $16$ combinations, and after simplifying the algebra we have the following. We denote
$F$ $=n^{3}n_{T}n_{C}(n_{T}-1)(n_{C}-1),$
$a_{T}$ $=\frac{n_{T}n_{C}-2n_{C}+n_{T}^{2}-2n_{T}+1}{n_{T}-1}$
$a_{C}$ $=\frac{n_{T}n_{C}-2n_{T}+n_{C}^{2}-2n_{C}+1}{n_{C}-1}$
$\overline{a}_{T}$ $=n_{C}n_{T}-3n_{C}+n_{T}^{2}-2n_{T}+1,$
$\overline{a}_{C}$ $=n_{C}n_{T}-3n_{T}+n_{C}^{2}-2n_{C}+1.$
Then
$T_{3}$ $=\frac{1}{F}\sum_{i\in[n]}\biggl[\sum_{k\in[n]:k\neq i}h_{\lambda ik}^{2}\frac{(n_{C}-1)(\widetilde{\bm{t}}^{(1)}_{k})^{2}+(n_{T}-1)(\widetilde{\bm{t}}^{(0)}_{k})^{2}}{\overline{h}_{i}^{2}}$ (31)
$-\frac{1}{n-2}\sum_{l\in[n]:l\neq i,k}\frac{h_{\lambda ik}h_{\lambda il}}{\overline{h}_{i}^{2}}((n_{C}-1)\widetilde{\bm{t}}^{(1)}_{k}\widetilde{\bm{t}}^{(1)}_{l}+(n_{T}-1)\widetilde{\bm{t}}^{(0)}_{k}\widetilde{\bm{t}}^{(0)}_{l})\biggr$] (32)
$+\frac{1}{F}\sum_{i,j\in[n]:i\neq j}\sum_{k\in[n]:k\neq i,j}\frac{h_{\lambda ik}h_{\lambda jk}}{\overline{h}_{i}\overline{h}_{j}}\Bigl[a_{T}(\widetilde{\bm{t}}^{(1)}_{k})^{2}-2n\widetilde{\bm{t}}^{(1)}_{k}\widetilde{\bm{t}}^{(0)}_{k}+a_{C}(\widetilde{\bm{t}}^{(0)}_{k})^{2}\Bigr]$ (33)
$+\frac{1}{n^{3}(n-1)}\sum_{i,j\in[n]:i\neq j}\biggl[\frac{h_{\lambda ij}^{2}}{\overline{h}_{i}\overline{h}_{j}}\Bigl[\frac{\widetilde{\bm{t}}^{(1)}_{i}\widetilde{\bm{t}}^{(1)}_{j}}{n_{T}(n_{T}-1)}+\frac{\widetilde{\bm{t}}^{(0)}_{i}\widetilde{\bm{t}}^{(0)}_{j}}{n_{C}(n_{C}-1)}+\frac{\widetilde{\bm{t}}^{(1)}_{i}\widetilde{\bm{t}}^{(0)}_{j}+\widetilde{\bm{t}}^{(0)}_{i}\widetilde{\bm{t}}^{(1)}_{j}}{n_{C}n_{T}}\Bigr]$ (34)
$-\frac{1}{n-2}\sum_{k\in[n]:k\neq i,j}\frac{h_{\lambda ji}}{\overline{h}_{i}\overline{h}_{j}}\Bigl[\frac{(h_{\lambda ik}\widetilde{\bm{t}}^{(1)}_{i}+h_{\lambda jk}\widetilde{\bm{t}}^{(1)}_{j})\widetilde{\bm{t}}^{(1)}_{k}}{n_{T}(n_{T}-1)}+\frac{(h_{\lambda ik}\widetilde{\bm{t}}^{(0)}_{i}+h_{\lambda jk}\widetilde{\bm{t}}^{(0)}_{j})\widetilde{\bm{t}}^{(0)}_{k}}{n_{C}(n_{C}-1)}$ (35)
$+\frac{(h_{\lambda ik}\widetilde{\bm{t}}^{(1)}_{i}+h_{\lambda jk}\widetilde{\bm{t}}^{(1)}_{j})\widetilde{\bm{t}}^{(0)}_{k}+(h_{\lambda ik}\widetilde{\bm{t}}^{(0)}_{i}+h_{\lambda jk}\widetilde{\bm{t}}^{(0)}_{j})\widetilde{\bm{t}}^{(1)}_{k}}{n_{C}n_{T}}\Bigr]$ (36)
[…]
To further simplify the above expression, note that it be written as the following quadratic form
$\begin{bmatrix}\widetilde{\bm{t}}^{(1)}\\ \widetilde{\bm{t}}^{(0)}\end{bmatrix}^{\top}\boldsymbol{Q}\begin{bmatrix}\widetilde{\bm{t}}^{(1)}\\ \widetilde{\bm{t}}^{(0)}\end{bmatrix}$,
where $\boldsymbol{Q}$ is a block matrix of the following form
$\boldsymbol{Q}=\begin{bmatrix}\boldsymbol{Q}^{00}&\boldsymbol{Q}^{01}\\ \boldsymbol{Q}^{10}&\boldsymbol{Q}^{11}\end{bmatrix}$,
and
$Q^{11}_{kk}$ $=\frac{n_{C}-1}{F}\sum_{i\in[n]:i\neq k}\frac{h_{\lambda ik}^{2}}{\overline{h}_{i}^{2}}+\frac{a_{T}}{F}\sum_{\begin{subarray}{c}i,j\in[n]:i\neq j\\ k\neq i,j\end{subarray}}\frac{h_{\lambda ik}h_{\lambda jk}}{\overline{h}_{i}\overline{h}_{j}},$
$Q^{00}_{kk}$ $=\frac{n_{T}-1}{F}\sum_{i\in[n]:i\neq k}\frac{h_{\lambda ik}^{2}}{\overline{h}_{i}^{2}}+\frac{a_{C}}{F}\sum_{\begin{subarray}{c}i,j\in[n]:i\neq j\\ k\neq i,j\end{subarray}}\frac{h_{\lambda ik}h_{\lambda jk}}{\overline{h}_{i}\overline{h}_{j}},$
$Q^{01}_{kk}=Q^{10}_{kk}$ $=-\frac{n}{F}\sum_{\begin{subarray}{c}i,j\in[n]:i\neq j\\ k\neq i,j\end{subarray}}\frac{h_{\lambda ik}h_{\lambda jk}}{\overline{h}_{i}\overline{h}_{j}},$
$Q^{11}_{k\ell}$ $=-\frac{n_{C}-1}{F(n-2)}\left(\sum_{i\in[n]:i\neq k,\ell}\frac{h_{\lambda ik}h_{\lambda i\ell}}{\overline{h}_{i}^{2}}\right)+\frac{1}{n^{3}(n-1)}\biggl[\frac{1}{n_{T}(n_{T}-1)}\frac{h_{\lambda k\ell}^{2}}{\overline{h}_{k}\overline{h}_{\ell}}$
$-\frac{1}{n-2}\frac{1}{n_{T}(n_{T}-1)}\biggl(\sum_{i\in[n]:i\neq k,\ell}\frac{h_{\lambda ki}h_{\lambda\ell k}}{\overline{h}_{k}\overline{h}_{i}}+\frac{h_{\lambda ik}h_{\lambda k\ell}}{\overline{h}_{i}\overline{h}_{\ell}}\biggr)$
$-\frac{1}{(n-2)n_{C}n_{T}(n-3)}\frac{\overline{a}_{T}}{n_{T}-1}\sum_{\begin{subarray}{c}i,j\in[n]:i\neq j\\ i,j\notin\{k,\ell\}\end{subarray}}\frac{h_{\lambda ik}h_{\lambda j\ell}}{\overline{h}_{i}\overline{h}_{j}}\biggr],$
$Q^{00}_{k\ell}$ $=-\frac{n_{T}-1}{F(n-2)}\left(\sum_{i\in[n]:i\neq k,\ell}\frac{h_{\lambda ik}h_{\lambda i\ell}}{\overline{h}_{i}^{2}}\right)+\frac{1}{n^{3}(n-1)}\biggl[\frac{1}{n_{C}(n_{C}-1)}\frac{h_{\lambda k\ell}^{2}}{\overline{h}_{k}\overline{h}_{\ell}}$
$-\frac{1}{n-2}\frac{1}{n_{C}(n_{C}-1)}\biggl(\sum_{i\in[n]:i\neq k,\ell}\frac{h_{\lambda ki}h_{\lambda\ell k}}{\overline{h}_{k}\overline{h}_{i}}+\frac{h_{\lambda ik}h_{\lambda k\ell}}{\overline{h}_{i}\overline{h}_{\ell}}\biggr)$
$-\frac{1}{(n-2)n_{C}n_{T}(n-3)}\frac{\overline{a}_{C}}{n_{C}-1}\sum_{\begin{subarray}{c}i,j\in[n]:i\neq j\\ i,j\notin\{k,\ell\}\end{subarray}}\frac{h_{\lambda ik}h_{\lambda j\ell}}{\overline{h}_{i}\overline{h}_{j}}\biggr],$
$Q^{01}_{k\ell}=Q^{10}_{\ell k}$ $=\frac{1}{n^{3}(n-1)}\biggl[\frac{1}{n_{C}n_{T}}\frac{h_{\lambda k\ell}^{2}}{\overline{h}_{k}\overline{h}_{\ell}}-\frac{1}{n-2}\frac{1}{n_{C}n_{T}}\biggl(\sum_{i\in[n]:i\neq k,\ell}\frac{h_{\lambda ki}h_{\lambda\ell k}}{\overline{h}_{k}\overline{h}_{i}}+\frac{h_{\lambda ik}h_{\lambda k\ell}}{\overline{h}_{i}\overline{h}_{\ell}}\biggr)$
$+\frac{n+1}{(n-2)n_{C}n_{T}(n-3)}\sum_{\begin{subarray}{c}i,j\in[n]:i\neq j\\ i,j\notin\{k,\ell\}\end{subarray}}\frac{h_{\lambda ik}h_{\lambda j\ell}}{\overline{h}_{i}\overline{h}_{j}}\biggr].$
Proof of Lemma˜6.
We argue in three steps.
Step 1: Existence and boundedness of $\boldsymbol{\beta}^{*}_{\lambda}$.
The first-order condition for the minimizer in (16) implies
$\boldsymbol{\beta}^{*}_{\lambda}\;=\;\lim_{n\to\infty}({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{X}}^{\top}\frac{\widetilde{{\boldsymbol{\mu}}}}{n}.$
By Assumption˜2, $n^{-1}{\bm{X}}^{\top}{\bm{X}}\to\Sigma_{X}$, where $\Sigma_{X}$ is bounded and invertible. Therefore
$\big\|({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}\big\|_{2}\;=\;O(n^{-1}).$ (38)
Moreover,
${\bm{X}}^{\top}\frac{\widetilde{{\boldsymbol{\mu}}}}{n}=\frac{1}{n}\sum_{j=1}^{n}{\bm{x}}_{j}\,\Big(n_{C}y_{j}^{(1)}+n_{T}y_{j}^{(0)}\Big).$
By Assumptions˜2 and 3, the averages $n^{-1}\sum_{j}{\bm{x}}_{j}y_{j}^{(1)}$ and $n^{-1}\sum_{j}{\bm{x}}_{j}y_{j}^{(0)}$ converge to finite limits, and $n_{T}/n$ and $n_{C}/n$ are bounded away from 0 and $1$. Hence $\big\|{\bm{X}}^{\top}\frac{\widetilde{{\boldsymbol{\mu}}}}{n}\big\|_{2}=O(n)$. Combining this with (38) shows that
$({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{X}}^{\top}\frac{\widetilde{{\boldsymbol{\mu}}}}{n}$
is $O(1)$ and converges to a finite limit. Thus $\boldsymbol{\beta}^{*}_{\lambda}$ exists and is bounded.
Step 2: Convergence of an auxiliary full-sample ridge coefficient.
For each $i\in[n]$, define the auxiliary (full-sample) ridge coefficient
$\widetilde{\boldsymbol{\beta}}^{(-i)}_{\lambda}:=\arg\min_{{\bm{b}}\in\mathbb{R}^{d}}\left\|{\bm{X}}{\bm{b}}-\widetilde{{\bm{y}}}^{(-i)}\right\|_{2}^{2}+\lambda\|{\bm{b}}\|_{2}^{2}=({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{X}}^{\top}\widetilde{{\bm{y}}}^{(-i)}.$
Unlike $\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$, this uses all $n$ rows of $\bm{X}$} rather than dropping unit $i$. We show that $\widetilde{\boldsymbol{\beta}}^{(-i)}_{\lambda}\to\boldsymbol{\beta}^{*}_{\lambda}$ in probability (with probability taken over the random assignment).
Algebraic manipulation yields $\widetilde{{\bm{y}}}^{(-i)}=\frac{\widetilde{{\boldsymbol{\mu}}}}{n}+\frac{{\bm{v}}^{(-i)}{\bm{z}}\widetilde{{\bm{t}}}^{(-i)}}{n}$ and ${\mathbb{E}}\big[\widetilde{{\bm{y}}}^{(-i)}\big]=\frac{\widetilde{{\boldsymbol{\mu}}}}{n}$. We denote the deviation with
$\Delta^{(-i)}\;:=\;\widetilde{{\bm{y}}}^{(-i)}-\frac{\widetilde{{\boldsymbol{\mu}}}}{n}.$
Because assignment is complete random assignment (sampling without replacement), the dependence across units is weak: covariances between $d_{j}$ and $d_{k}$ are of order $1/n$ for $j\neq k$. Using this standard finite-population sampling algebra and the bounded fourth-moment assumptions in Assumptions˜1 and 2, one obtains
$\big\|{\bm{X}}^{\top}\Delta^{(-i)}\big\|_{2}=O(\sqrt{n}),$
uniformly in $i$. Intuitively, $\Delta^{(-i)}$ is a mean-zero perturbation with components of order $1$, and summing ${\bm{x}}_{j}\Delta^{(-i)}_{j}$ accumulates fluctuations of order $\sqrt{n}$ in Euclidean norm.
We may now write
${\bm{X}}^{\top}\widetilde{{\bm{y}}}^{(-i)}={\bm{X}}^{\top}\frac{\widetilde{{\boldsymbol{\mu}}}}{n}+{\bm{X}}^{\top}\Delta^{(-i)}.$
Premultiplying by $({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}$ gives
$\widetilde{\boldsymbol{\beta}}^{(-i)}_{\lambda}=({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{X}}^{\top}\frac{\widetilde{{\boldsymbol{\mu}}}}{n}\;+\;({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{X}}^{\top}\Delta^{(-i)}.$
The first term converges to $\boldsymbol{\beta}^{*}_{\lambda}$ by Step 1. For the second term, we use $\|({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}\|_{2}=O(n^{-1})$ and $\|{\bm{X}}^{\top}\Delta^{(-i)}\|_{2}=O(\sqrt{n})$, which implies
$\left\|({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{X}}^{\top}\Delta^{(-i)}\right\|_{2}=O(n^{-1})\cdot O(\sqrt{n})=O(n^{-1/2})\;\to\;0.$
Hence
$\widetilde{\boldsymbol{\beta}}^{(-i)}_{\lambda}\;\xrightarrow{p}\;\boldsymbol{\beta}^{*}_{\lambda},\qquad\text{for each }i.$
Step 3: Leave-one-out versus full-sample ridge.
We now compare $\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$, which regresses $\widetilde{{\bm{y}}}^{(-i)}_{-i}$ on ${\bm{X}}_{-i}$, to $\widetilde{\boldsymbol{\beta}}^{(-i)}_{\lambda}$, which regresses $\widetilde{{\bm{y}}}^{(-i)}$ on $\bm{X}$}. The standard leave-one-out identity for ridge regression (cf. Lemma˜B.1) yields
$\widetilde{\boldsymbol{\beta}}^{(-i)}_{\lambda}-\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}=\frac{({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{x}}_{i}\left(\widetilde{y}^{(-i)}_{i}-{\bm{x}}_{i}^{\top}\widetilde{\boldsymbol{\beta}}^{(-i)}_{\lambda}\right)}{1-h_{\lambda ii}},$ (39)
where
$h_{\lambda ii}={\bm{x}}_{i}^{\top}({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{x}}_{i}$
is the ridge leverage score of unit $i$. By Assumption˜2, $\|{\bm{x}}_{i}\|_{2}$ is uniformly bounded. Using (38),
$h_{\lambda ii}\;\leq\;\|{\bm{x}}_{i}\|_{2}^{2}\,\big\|({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}\big\|_{2}=O(n^{-1}),$
so $1-h_{\lambda ii}\to 1$ uniformly in $i$. Moreover, $\|({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{x}}_{i}\|_{2}=O(n^{-1})$, and $\big|\widetilde{y}^{(-i)}_{i}-{\bm{x}}_{i}^{\top}\widetilde{\boldsymbol{\beta}}^{(-i)}_{\lambda}\big|=O(1)$, because $\widetilde{y}^{(-i)}_{i}$ is a reweighting of $y_{i}^{(1)}$ or $y_{i}^{(0)}$ with coefficients of order $1$, and $\widetilde{\boldsymbol{\beta}}^{(-i)}_{\lambda}$ is bounded by the conclusion in Step 2. Thus the right-hand side of (39) is $O(n^{-1})$, uniformly in $i$, and therefore
$\big\|\widetilde{\boldsymbol{\beta}}^{(-i)}_{\lambda}-\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}\big\|_{2}\;\to\;0.$
Finally, Step 2 shows $\widetilde{\boldsymbol{\beta}}^{(-i)}_{\lambda}\xrightarrow{p}\boldsymbol{\beta}^{*}_{\lambda}$ and Step 3 shows $\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}-\widetilde{\boldsymbol{\beta}}^{(-i)}_{\lambda}\to\boldsymbol{0}$. Combining these, for each $i$,
$\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}\;\xrightarrow{p}\;\boldsymbol{\beta}^{*}_{\lambda}.$
Proof of Theorem˜7.
Recall that
$\widehat{\tau}_{\textup{LDM}}=\frac{1}{n_{T}}\sum_{i:d_{i}=1}\big(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}\big)\;-\;\frac{1}{n_{C}}\sum_{i:d_{i}=0}\big(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}\big)$,
where $d_{i}\in\{0,1\}$ is the treatment indicator, $n_{T}=\sum_{i}d_{i}$, $n_{C}=n-n_{T}$, and $\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$ is the leave-one-out ridge coefficient for unit $i$.
By Lemma˜6,
$\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}\;\xrightarrow{p}\;\boldsymbol{\beta}^{*}_{\lambda}\quad\text{for each }i.$
Define the infeasible “oracle” regression-adjusted difference-in-means estimator that uses $\boldsymbol{\beta}^{*}_{\lambda}$ in place of $\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$:
$\widehat{\tau}^{*}_{\textup{LDM}}:=\frac{1}{n_{T}}\sum_{i:d_{i}=1}\big(y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}^{*}_{\lambda}\big)\;-\;\frac{1}{n_{C}}\sum_{i:d_{i}=0}\big(y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}^{*}_{\lambda}\big).$
A standard Slutsky-type argument (using linearity of averages and the fact that $n_{T},n_{C}$ are of order $n$) implies
$\widehat{\tau}_{\textup{LDM}}-\widehat{\tau}^{*}_{\textup{LDM}}\;\xrightarrow{p}\;0.$
Therefore,
$\sqrt{n}\,(\widehat{\tau}_{\textup{LDM}}-\tau)\quad\text{and}\quad\sqrt{n}\,(\widehat{\tau}^{*}_{\textup{LDM}}-\tau)$
have the same asymptotic distribution. It remains to analyze $\widehat{\tau}^{*}_{\textup{LDM}}$.
Now note that since $\boldsymbol{\beta}_{\lambda}^{*}$ is a fixed vector, by (13), the limiting variance of $\widehat{\tau}^{*}_{\textup{LDM}}$ is given by
$\lim_{n\to\infty}\frac{1}{n}\big\|({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}^{*}_{\lambda}-({\boldsymbol{\mu}}-\overline{{\boldsymbol{\mu}}})\big\|_{2}^{2}.$
Because complete random assignment corresponds to sampling $n_{T}$ treated units without replacement from a finite population of fixed potential outcomes and covariates, and because $n_{T}/n$ and $n_{C}/n$ are bounded away from 0 and $1$ by Assumption˜3, Hájek’s finite-population central limit theorem for sampling without replacement (hajek1960limiting) applies to $\sqrt{n}\,(\widehat{\tau}^{*}_{\textup{LDM}}-\tau)$, under our bounded fourth-moment conditions (Assumptions˜1 and 2). In particular,
$\sqrt{n}\,(\widehat{\tau}^{*}_{\textup{LDM}}-\tau)\;\xrightarrow{d}\;\mathcal{N}\!\Big(0,\;\lim_{n\to\infty}\frac{1}{n}\big\|({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}^{*}_{\lambda}-({\boldsymbol{\mu}}-\overline{{\boldsymbol{\mu}}})\big\|_{2}^{2}\Big).$
Finally, since $\widehat{\tau}_{\textup{LDM}}-\widehat{\tau}^{*}_{\textup{LDM}}\;\xrightarrow{p}\;0$, Slutsky’s theorem implies that the same asymptotic normal limit holds for $\sqrt{n}(\widehat{\tau}_{\textup{LDM}}-\tau)$. ∎